### v1

# Scalable and robust regression models for continuous proportional data

Abstract
Beta regression is used routinely for continuous proportional data, but it often encounters practical issues such as a lack of robustness of regression parameter estimates to misspecification of the beta distribution. We develop an improved class of generalized linear models starting with the continuous binomial (cobin) distribution and further extending to dispersion mixtures of cobin distributions (micobin). The proposed cobin regression and micobin regression models have attractive robustness, computation, and flexibility properties. A key innovation is the Kolmogorov-Gamma data augmentation scheme, which facilitates Gibbs sampling for Bayesian computation, including in hierarchical cases involving nested, longitudinal, or spatial data. We demonstrate robustness, ability to handle responses exactly at the boundary (0 or 1), and computational efficiency relative to beta regression in simulation experiments and through analysis of the benthic macroinvertebrate multimetric index of US lakes using lake watershed covariates.
Keywords: Bayesian; Data augmentation; Generalized linear model; Latent Gaussian model; Markov chain Monte Carlo.
1 Introduction
Regression analysis of proportional data, corresponding to bounded continuous data in the unit interval $[0,1]$, is a common focus in many fields. Such data arise in diverse contexts, ranging from the analysis of rates or indices in econometrics (Papke and Wooldridge,, 1996) to the percentage of tissue area in medical imaging (Peplonska et al.,, 2012). We are particularly motivated by ecological applications; Warton and Hui, (2011) estimates $\sim 14\%$ of ecology articles involve data based on proportions that are not derived from counts, with examples including measurements of percent coverage (Korhonen et al.,, 2024) and ecological indices (Lindholm et al.,, 2021).
Generalized linear models (GLMs) (Nelder and Wedderburn,, 1972) and their variants are fundamental tools for regression analysis. GLMs have multiple appealing theoretical and computational properties (McCullagh and Nelder,, 1989). For continuous proportional data, one of two strategies is typically taken. Firstly, one may transform the data to fall on the real line and then apply a Gaussian linear model. However, such transformation-based approaches complicate the interpretation of the results in the original scale and have problems when some observations are concentrated near the boundary of the support. Alternatively, beta regression is widely applied (Ferrari and Cribari-Neto,, 2004; Cribari-Neto and Zeileis,, 2010), assuming beta-distributed response variables supported on the open interval (0,1). We refer to Douma and Weedon, (2019) for a review of beta regression and applications in ecological contexts.
However, beta regression has several limitations. First, beta distributions are not in the natural exponential family (Morris,, 1982) and thus do not strictly belong to the GLM class. This implies that the mean and dispersion parameters are not orthogonal to each other (Ferrari and Cribari-Neto,, 2004), and the well-established properties of GLMs may not hold. Beta regression also faces computational difficulties when considering extensions to accommodate complex dependence structures. This limitation is particularly relevant in Bayesian frameworks, which support flexible hierarchical extensions (Bolker et al.,, 2009). Finally, the presence of exact 0s or 1s prevents the direct application of beta regression models. This issue is often bypassed by manipulating the data to be within the open interval (0,1) (Smithson and Verkuilen,, 2006), but the results are often highly sensitive to such preprocessing (Kosmidis and Zeileis,, 2024).
The motivation of this article is to introduce a proper GLM approach to continuous proportional response data, without the need for data preprocessing and facilitating computation, including in complex settings involving random effects. With these goals in mind, we propose continuous binomial (cobin) regression, a name inspired by models based on the continuous Bernoulli distribution (Loaiza-Ganem and Cunningham,, 2019; Quijano Xacur,, 2019). The cobin distribution is an exponential dispersion model (Jørgensen,, 1987) with orthogonal mean and dispersion parameters, implying that the corresponding GLM has appealing properties, including consistency of maximum likelihood estimates (MLE) under model misspecification. For posterior sampling using the canonical link function, we propose a novel data augmentation strategy based on Kolmogorov-Gamma random variables, leading to a conditionally normal likelihood, similar to Pólya-Gamma augmentation for logistic models (Polson et al.,, 2013). We develop an efficient sampler for Kolmogorov-Gamma random variables. The proposed Gibbs sampler can easily accommodate extensions to Gaussian latent variables and random effects, including in spatial contexts. Furthermore, we show uniform ergodicity of the proposed MCMC algorithms, providing theoretical guarantees on efficient inference procedure.
We also propose micobin regression, an extension of cobin regression based on dispersion mixtures of continuous binomial (micobin) distributions. This improves flexibility and robustness by localizing the dispersion parameter (Wang and Blei,, 2018), and can handle exact 0s or 1s without modifying the data. Micobin regression addresses non-structural boundary values (Blasco-Moreno et al.,, 2019) and is distinct from models that explicitly handle boundary values by assigning positive probability mass to 0s or 1s, such as zero/one inflated or censored models (Ospina and Ferrari,, 2012; Kubinec,, 2023; Kosmidis and Zeileis,, 2024). We describe how micobin regression can be further extended to allow the dispersion parameter to systematically vary according to covariates. Through simulations, we show that micobin regression achieves better predictive performance than beta regression in misspecified settings.
As an illustration, we analyze the benthic macroinvertebrate multimetric index (MMI) of US lakes, an index ranging from 0 to 100 (scaled by 0.01 throughout the article) that reflects the condition of lake macroinvertebrate communities (Stoddard et al.,, 2008). We examine association between MMI and lake watershed covariates, and use them to predict MMI at unsampled lakes; see Figure 1. Since accounting for spatial dependence is highly important in such spatial ecological data (Guélat and Kéry,, 2018), we fit spatial cobin and micobin regression models with latent Gaussian process random effects. We compare the results with those from spatial beta regression, highlighting the robustness and computational advantages provided by Kolmogorov-Gamma augmentation.
Figure 1: (Left) Benthic macroinvertebrate multimetric index of 949 lakes in conterminous US (U.S. Environmental Protection Agency,, 2022) (Right) Urban land cover in the watersheds of 55,215 lakes, the log-transformed percentage of watershed area classified as developed, medium and high intensity land use (2016 National Land Cover Database classes 23, 24) (Hill et al.,, 2018).
The paper is structured as follows. In Section 2, we introduce cobin and micobin regressions, study their properties, and present hierarchical extensions. In Section 3, we establish the Kolmogorov-Gamma integral identity in Theorem 1, which could be of independent interest for other models, and describe our Kolmogorov-Gamma augmentation strategy for posterior computation. In Section 4, we develop a highly efficient algorithm to sample Kolmogorov-Gamma random variables. In Section 5 and 6, we perform simulation studies to support our claims and present an application example. All proofs of theorems are in Appendix S.1 except for Theorem 1.
2 Cobin and micobin regression models
2.1 Continuous binomial distribution
An exponential dispersion model (Jørgensen,, 1987) is an extension of a one-parameter natural exponential family by adding a dispersion parameter that controls variance in a systematic way. Notable examples of such extensions include normal with fixed variance to normal distributions with unknown variance, exponential to gamma distributions, and geometric to negative binomial distributions. We refer to Jørgensen, (1992) for a comprehensive review of exponential dispersion models and their properties.
Our goal is to find a family of distributions supported on the unit interval that serve as the random component of a GLM having appealing properties. Unlike binary regression where the conditional mean completely determines the distribution, it is desirable to have an additional dispersion parameter instead of just the natural parameter. We introduce the continuous binomial (cobin) distribution, which is in the exponential dispersion family and contains the uniform distribution as a special case. We first define the density of the cobin distribution. To avoid overloading notation, in what follows, we assume that $(e^{x}-1)/x$, $\sinh(x)/x$, and similar expressions are well defined in the limit at $x=0$.
Definition 1 (cobin).
The continuous binomial (cobin) distribution with natural parameter $\theta\in\mathbb{R}$ and dispersion parameter $\lambda^{-1}\in\{1,1/2,1/3,\dots\}$, denoted as $Y\sim\mathrm{cobin}(\theta,\lambda^{-1})$, is a exponential dispersion model with density function
$p_{\mathrm{cobin}}(y;\theta,\lambda^{-1})=h(y,\lambda)\exp\left[\lambda\{\theta y-B(\theta)\}\right]=h(y,\lambda)\frac{e^{\lambda\theta y}}{\{(e^{\theta}-1)/\theta\}^{\lambda}},\quad 0\leq y\leq 1,$ (1)
where $B(\theta)=\log\left\{(e^{\theta}-1)/\theta\right\}$ and $h(y,\lambda)=\frac{\lambda}{(\lambda-1)!}\sum_{k=0}^{\lambda}(-1)^{k}{\lambda\choose k}\{\max(\lambda y-k,0)\}^{\lambda-1}$. The support is $[0,1]$ when $\lambda=1$ and $(0,1)$ if $\lambda\geq 2$.
When $\theta=0$ and $\lambda=1$, the cobin distribution reduces to $\mathrm{Unif}(0,1)$. In Appendix S.3, we provide a formal derivation of the cobin as an exponential dispersion model, including a proof that $\lambda$ must be an integer, along with comparable derivations for the normal, gamma, and inverse Gaussian families. The cobin distribution has appeared in the literature in various contexts, dating back to Bates, (1955) in the study of time intervals between accidents. The special case of $\lambda=1$, with different parameterizations based on $\varphi=1/(1+e^{-\theta})$, is called the continuous Bernoulli distribution (Loaiza-Ganem and Cunningham,, 2019), since the density is proportional to $\varphi^{y}(1-\varphi)^{1-y}$. The name “continuous binomial” comes from the fact that the cobin distribution arises from the convolution of i.i.d. continuous Bernoulli random variables, $Y_{1},\dots,Y_{\lambda}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}\mathrm{cobin}(\theta,1)$ then $\lambda^{-1}\sum_{l=1}^{\lambda}Y_{l}\sim\mathrm{cobin}(\theta,\lambda^{-1})$, which follows from the convolution property of the exponential dispersion model (Jørgensen,, 1987).
By the properties of the exponential family, the mean and variance can be derived directly by differentiating $B(\theta)=\log\left\{(e^{\theta}-1)/\theta\right\}$,
$E(Y)=B^{\prime}(\theta)=e^{\theta}/(e^{\theta}-1)-\theta^{-1},\quad\operatorname*{var}(Y)=\lambda^{-1}B^{\prime\prime}(\theta)=\lambda^{-1}\{(2-e^{\theta}-e^{-\theta})^{-1}+\theta^{-2}\},$
which illustrates that $E(Y)$ is only controlled by the natural parameter $\theta$ and $\operatorname*{var}(Y)$ is proportional to the dispersion parameter $\lambda^{-1}$. See Figure 2 for examples and a comparison with the beta distribution having the same mean and variance.
An important distinction from the beta distribution is that the cobin distribution does not allow bimodal densities having spikes at both zero and one. This implies that the range of possible variances under the cobin distribution is smaller than that for that beta, and the maximum variance $1/12$ is achieved at $\mathrm{Unif}(0,1)$. This restriction may be an advantage in many applications in which the true density of the response given covariates is unlikely to be U-shaped, but U-shaped fitted distributions can arise in beta regression for certain covariate values due to lack of fit and sparse data.
Figure 2: Comparison of beta, cobin, and mixture of cobin (micobin) with common mean and variance.
2.2 Cobin regression
With a link function $g:(0,1)\to\mathbb{R}$ that is strictly monotone and differentiable, the GLM with continuous binomial response can be expressed as:
$Y_{i}\mid\theta_{i},\lambda\stackrel{{\scriptstyle\mathrm{ind}}}{{\sim}}\mathrm{cobin}(\theta_{i},\lambda^{-1}),\quad\theta_{i}=(B^{\prime})^{-1}\{g^{-1}(\bm{x}_{i}^{\mathrm{\scriptscriptstyle T}}\bm{\beta})\},\quad i=1,\dots,n,$ (2)
which implies $g\{E(y_{i}\mid x_{i})\}=\bm{x}_{i}^{\mathrm{\scriptscriptstyle T}}\bm{\beta}$. We refer to the model (2) as cobin regression. Denoting the linear predictor $\eta_{i}=\bm{x}_{i}^{\mathrm{\scriptscriptstyle T}}\bm{\beta}$ and the conditional mean $\mu_{i}=g^{-1}(\eta_{i})$, the likelihood equations for $\bm{\beta}$ are
$\frac{\partial}{\partial\beta_{j}}\sum_{i=1}^{n}\log p_{\mathrm{cobin}}(y_{i};\theta_{i},\lambda^{-1})=\lambda\sum_{i=1}^{n}\frac{(y_{i}-\mu_{i})x_{ij}}{B^{\prime\prime}(\theta_{i})}\frac{\partial\mu_{i}}{\partial\eta_{i}}=0,\quad j=1,\dots,p$ (3)
where the solution $\hat{\bm{\beta}}$ does not depend on the dispersion parameter $\lambda^{-1}$. Standard procedures, such as Newton–Raphson or iteratively reweighted least squares, can be used to find the MLE as well as maximum a posteriori (MAP) estimates.
Unlike beta regression, where two shape parameters corresponding to sufficient statistics $T(y)=\{\log(y),\log(1-y)\}$ jointly control the mean, cobin regression is a natural exponential family with a natural parameter $\theta$ corresponding to sufficient statistics $T(y)=y$. Therefore, unlike beta regression, cobin regression is a proper GLM and hence inherits many attractive properties, such as a concave log-likelihood function under the canonical link. We highlight the robustness property of the solution of (3) against model misspecification.
Proposition 1.
Under mild regularity conditions (Gourieroux et al.,, 1984), $\hat{\bm{\beta}}$ of the cobin regression likelihood equations (3) is consistent if $E(y_{i}\mid x_{i})=g^{-1}(\bm{x}_{i}^{\mathrm{\scriptscriptstyle T}}\bm{\beta})$ is correctly specified.
Since the cobin distribution with fixed $\lambda$ belongs to the natural exponential family, Proposition 1 is a direct consequence of viewing (3) as quasi-maximum likelihood estimating equations associated with a natural exponential family (Gourieroux et al.,, 1984); also see §4.2.6 and §8.3 of Agresti, (2015) for details. Thus, under the correct mean structure, the cobin regression produces an asymptotically valid point estimate even if $Y$ does not follow the cobin distribution.
Another important difference between beta and cobin is the form of the score function, the derivative of the log-likelihood with respect to $\bm{\beta}$. In contrast to the cobin score function (3) which is bounded in terms of $y$, the beta score function depends on $\log\{y/(1-y)\}$ and is thus unbounded for $y\in(0,1)$ (Ferrari and Cribari-Neto,, 2004). Thus, $y$ near its boundary values can lead to an arbitrarily large change in the beta regression log-likelihood landscape. Given the fact that boundedness of the score function often serves as a necessary condition for robustness properties (e.g., Prop. 2 of Cantoni and Ronchetti, (2001)), cobin regression has greater robustness to observations near the boundaries than beta regression.
The canonical link function of cobin regression, which we call the cobit link, satisfies $g_{\mathrm{cobit}}^{-1}(\eta)=B^{\prime}(\eta)=e^{\eta}/(e^{\eta}-1)-\eta^{-1}$. Although $g_{\mathrm{cobit}}(\mu)$ does not have a closed form expression, it can be easily inverted numerically in practice. Compared to logistic function $1/(1+e^{-\eta})$, $g_{\mathrm{cobit}}^{-1}(\eta)$ slowly approaches 0 and 1 as $|\eta|\to\infty$, asymptotically at the same speed as the Cauchy distribution c.d.f. See Figure S.5 in Appendix S.3 for details. Similarly to heavy-tailed link functions in binary response regression, this behavior is appealing in allowing the mean parameter to approach zero or one more slowly for extreme values of the predictors, reducing sensitivity to certain types of outliers.
Under the canonical link function, the cobin regression model can be expressed as $Y_{i}\mid\bm{x}_{i}\stackrel{{\scriptstyle\mathrm{ind}}}{{\sim}}\mathrm{cobin}(\bm{x}_{i}^{\mathrm{\scriptscriptstyle T}}\bm{\beta},\lambda^{-1})$ for $i=1,\dots,n$. In the following sections, we focus on cobin regression with canonical link functions, which greatly simplifies model fitting. Regardless of the choice of the link function, relationships between the conditional mean and covariates can be reported in terms of marginal effects (Williams,, 2012).
2.3 Mixture of cobin regression and varying dispersion model
Cobin regression comes with some limitations. First, the parameter $\lambda$ is only allowed to be an integer. This not only limits flexibility, but also introduces significant difficulties when we want to allow the dispersion parameter to vary systematically based on covariates. Next, the support of the cobin distribution is an open interval $(0,1)$ when $\lambda\geq 2$ and therefore cannot handle data at exact boundaries. Beta regression suffers from the same problem, and while “nudging” data into $(0,1)$ before analysis (Smithson and Verkuilen,, 2006) is a standard practice in the literature, doing so is certainly not desirable.
To overcome these limitations, we introduce an extension of cobin regression, named micobin regression. Specifically, we propose using the cobin distribution dispersion mixture as a response distribution, defined as follows.
Definition 2 (micobin).
We say $Y$ follows a mixture of cobin distributions (micobin) with natural parameter $\theta\in\mathbb{R}$ and dispersion parameter $\psi\in(0,1)$, written as $Y\sim\mathrm{micobin}(\theta,\psi)$, if
$Y\mid\lambda\sim\mathrm{cobin}(\theta,\lambda^{-1}),\quad(\lambda-1)\sim\mathrm{negbin}(2,\psi),$ (4)
where $\mathrm{negbin}(r,\psi)$ denotes a negative binomial with mean $r(1-\psi)/\psi$ and variance $r(1-\psi)/\psi^{2}$. The mean and variance are $E(Y)=B^{\prime}(\theta)$ and $\operatorname*{var}(Y)=\psi B^{\prime\prime}(\theta)$.
The micobin distribution is obtained by mixing the cobin distributions over the dispersion parameter $\lambda$, analogous to obtaining Student’s $t$ distribution as a scale mixture of normal distributions. See Figure 2 for comparisons with the cobin and beta distributions ($\psi=1$ corresponds to the limiting case). The mixture over the dispersion parameter preserves the mean structure and improves the robustness against outliers (Wang and Blei,, 2018). The choice of $(\lambda-1)\sim\mathrm{negbin}(2,\psi)$ leads to a property that $E(\lambda^{-1})=\psi$, which implies $\operatorname*{var}(Y)=\psi B^{\prime\prime}(\theta)$. Thus, $\psi$ plays a similar role as $\lambda^{-1}$ as a dispersion parameter, with the additional flexibility that $\psi$ can take any value between 0 and 1.
The micobin distribution is supported on the closed interval $[0,1]$ for any choice of $\psi\in(0,1)$, with positive densities at the boundaries $p_{\mathrm{micobin}}(0;\theta,\psi)=\psi^{2}\theta/(e^{\theta}-1)>0$ and $p_{\mathrm{micobin}}(1;\theta,\psi)=\psi^{2}\theta e^{\theta}/(e^{\theta}-1)>0$. Thus, it accommodates boundary data without needing to arbitrarily nudge boundary values to lie in the open interval $(0,1)$. Micobin differs from models that assign positive point mass at 0 and/or 1, which treat boundary values as qualitatively different from others; we refer to Blasco-Moreno et al., (2019) for a detailed discussion of structural and non-structural boundary values, focusing on ecological contexts.
Under the canonical link, the micobin regression model can be compactly written as
$y_{i}\mid\bm{x}_{i}\stackrel{{\scriptstyle\mathrm{ind}}}{{\sim}}\mathrm{micobin}(\bm{x}_{i}^{\mathrm{\scriptscriptstyle T}}\bm{\beta},\psi),\quad i=1,\dots,n,$ (5)
which implies $g_{\mathrm{cobit}}\{E(y_{i}\mid x_{i})\}=\bm{x}_{i}^{\mathrm{\scriptscriptstyle T}}\bm{\beta}$. Micobin regression can be extended to allow the dispersion to vary systematically with covariates, which may differ from covariates in the mean regression component of the model (Smyth,, 1989; Smithson and Verkuilen,, 2006). The varying dispersion micobin regression, under the canonical link for mean and logit link for dispersion, can be written as
$Y_{i}\mid x\bm{}i,\psi_{i}\stackrel{{\scriptstyle\mathrm{ind}}}{{\sim}}\mathrm{micobin}(\bm{x}_{i}^{\mathrm{\scriptscriptstyle T}}\bm{\beta},\psi_{i}),\quad\mathrm{logit}(\psi_{i})=\bm{d}_{i}^{\mathrm{\scriptscriptstyle T}}\bm{\gamma},\quad i=1,\dots,n,$ (6)
where $\bm{d}_{i}\in\mathbb{R}^{d}$ is a dispersion covariate that may overlap with $\bm{x}_{i}$, and $\bm{\gamma}\in\mathbb{R}^{d}$ is the coefficient.
2.4 Extensions to mixed effects models with latent Gaussian structure
When data comes with multilevel, longitudinal, or spatial structure, cobin and micobin regression models can be naturally extended to mixed-effect models, and more generally to latent Gaussian models (Rue et al.,, 2009). For spatially indexed proportional data $y(s)$ and covariate $\bm{x}(s)$, we focus on spatial generalized linear mixed models (Diggle et al.,, 1998) with cobit link,
$g_{\mathrm{cobit}}(E\{y(s)\mid\bm{x}(s),u(s)\})=\bm{x}(s)^{\mathrm{\scriptscriptstyle T}}\bm{\beta}+u(s),\quad u(\cdot)\sim\mathrm{mean\,\,zero\,\,Gaussian\,\,process}.$ (7)
With a choice of response distribution such as cobin or micobin, this allows us to perform both the inference of $\bm{\beta}$ and prediction of response at new locations in a single modeling framework (Banerjee et al.,, 2014).
Inference for non-Gaussian spatial models (7) is often performed using Bayesian approaches implemented with MCMC, due to the challenges in inferring spatial random effects at many locations. However, generic sampling algorithms such as Metropolis-Hastings require careful tuning and face significant challenges when sampling high-dimensional latent parameters. In the following section, we propose a novel data augmentation scheme called Kolmogorov-Gamma augmentation, which converts cobin or micobin likelihoods into a conditionally normal likelihood and yields a simple Gibbs sampler that does not require tuning and with theoretical guarantees on rapid convergence.
3 Inference with Kolmogorov-Gamma augmentation
3.1 Kolmogorov-Gamma distribution and integral identity
First, we define Kolmogorov-Gamma random variables.
Definition 3.
We say a positive random variable $\kappa$ follows a Kolmogorov-Gamma (KG) distribution with parameters $b>0$ and $c\in\mathbb{R}$, denoted as $\kappa\sim\mathrm{KG}(b,c)$, if
$\kappa\stackrel{{\scriptstyle\mathrm{d}}}{{=}}\frac{1}{2\pi^{2}}\sum_{k=1}^{\infty}\frac{\epsilon_{k}}{k^{2}+c^{2}/(4\pi^{2})},\quad\epsilon_{k}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}\mathrm{Gamma}(b,1),\quad k=1,2,\dots,$ (8)
where =$\stackrel{{\scriptstyle\mathrm{d}}}{{=}}$ denotes equality in distribution.
For the case where $b=1$ and $c=0$, $\mathrm{KG}(1,0)$ scaled by $\pi^{2}$ corresponds to the squared Kolmogorov (or squared Kolmogorov-Smirnov) distribution, the infinite convolution of independent exponential distributions (Andrews and Mallows,, 1974, §4). Since $\epsilon_{k}$ are gamma distributed, following a similar naming convention as Pólya-Gamma (Polson et al.,, 2013), we call $\kappa$ a Kolmogorov-Gamma random variable. The difference with the Pólya-Gamma is the term $k^{2}$ in the denominator in (8), which is $(k-0.5)^{2}$ for Pólya-Gamma. From relationships between exponential and gamma distributions, $\mathrm{KG}(b,c)$ with integer $b$ is equal in distribution to the sum of $b$ independent $\mathrm{KG}(1,c)$ variables. In Section 4, we study the density and random-variate generating scheme in depth.
Now we describe the key result. Under the canonical link, the cobin likelihood (1) is proportional to $(e^{\eta})^{\lambda y}/\{(e^{\eta}-1)/\eta\}^{\lambda}$ in terms of the linear predictor $\eta=\theta=\bm{x}^{\mathrm{\scriptscriptstyle T}}\bm{\beta}$, which is not a well-recognized form in terms of $\eta$. We present the following Kolmogorov-Gamma integral identity, which establishes a direct connection between the cobin likelihood and Kolmogorov-Gamma random variables.
Theorem 1.
For any $a\in\mathbb{R}$ and $b>0$, the following integral identity holds for $\eta\in\mathbb{R}$:
$\frac{(e^{\eta})^{a}}{\{(e^{\eta}-1)/\eta\}^{b}}=e^{(a-b/2)\eta}\int_{0}^{\infty}e^{-\kappa\eta^{2}/2}{p_{\textsc{kg}}}(\kappa;b,0)\mathrm{d}\kappa,$ (9)
where ${p_{\textsc{kg}}}(\kappa;b,0)$ is the density of a $\mathrm{KG}(b,0)$ random variable.
Proof.
It suffices to show the result when $a=0$, since $e^{a\eta}$ is always positive and cancels out. Using the fact that the Laplace transformation of $\epsilon_{k}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}\mathrm{Gamma}(b,1)$ is $E\{\exp(-\epsilon_{k}t)\}=(1+t)^{-b}$, the Laplace transformation of $\kappa\sim\mathrm{KG}(b,0)$ is
$E\{\exp(-\kappa t)\}=\prod_{k=1}^{\infty}E\left\{\exp\left(-\frac{\epsilon_{k}t}{2\pi^{2}k^{2}}\right)\right\}=\prod_{k=1}^{\infty}\left(1+\frac{t}{2\pi^{2}k^{2}}\right)^{-b}=\left[\frac{(t/2)^{1/2}}{\sinh\{(t/2)^{1/2}\}}\right]^{b},$ (10)
where the last equation follows from the Weierstrass factorization theorem (Olver et al.,, 2010, §4.36.1). Plugging in $t=\eta^{2}/2$, we have $E(e^{-\eta^{2}\kappa/2})=\{(\eta/2)/\sinh(\eta/2)\}^{b}=e^{b\eta/2}/\{(e^{\eta}-1)/\eta\}^{b}$, which completes the proof. ∎
Theorem 2 shows that the conditional distribution with density $p(\kappa\mid\eta)\propto e^{-\kappa\eta^{2}/2}{p_{\textsc{kg}}}(\kappa;b,0)$ is $\mathrm{KG}(b,\eta)$, which arises from treating the integrand of (9) as an unnormalized density of $\kappa$.
Theorem 2.
The Kolmogorov-Gamma random variable $\mathrm{KG}(b,c)$ in Definition 3 coincides with the distribution arising from exponential tilting of the $\mathrm{KG}(b,0)$ random variable, with density equal to ${p_{\textsc{kg}}}(x;b,c)=\left\{\sinh(c/2)/(c/2)\right\}^{b}\exp(-c^{2}x/2){p_{\textsc{kg}}}(x;b,0)$, $x>0$.
3.2 Kolmogorov-Gamma augmentation and blocked Gibbs sampler
Based on Theorem 1, consider the augmented model by introducing KG variables $\bm{\kappa}=\{\kappa_{i}\}_{i=1}^{n}$,
$p(y_{i},\kappa_{i}\mid\bm{x}_{i},\bm{\beta},\lambda)=h(y_{i},\lambda)\exp\{{\lambda(y_{i}-0.5)\bm{x}_{i}^{\mathrm{\scriptscriptstyle T}}\bm{\beta}}-\kappa_{i}(\bm{x}_{i}^{\mathrm{\scriptscriptstyle T}}\bm{\beta})^{2}/2\}{p_{\textsc{kg}}}(\kappa_{i};\lambda,0),$ (11)
for $i=1,\dots,n$, which reduces to original cobin regression upon marginalization $\int_{0}^{\infty}p(y_{i},\kappa_{i}\mid x_{i},\bm{\beta},\lambda)\mathrm{d}\kappa_{i}=p_{\mathrm{cobin}}(y_{i};\bm{x}_{i}^{\mathrm{\scriptscriptstyle T}}\bm{\beta},\lambda^{-1})$. Then, the log-likelihood from (11) conditional on $\bm{\kappa}$ and $\lambda$ becomes a quadratic form in terms of $\bm{\beta}$. In addition, the conditional distribution of $\kappa_{i}$ given $\bm{\beta}$ and $\lambda$ is $\mathrm{KG}(\lambda,\bm{x}_{i}^{\mathrm{\scriptscriptstyle T}}\bm{\beta})$ according to Theorem 2. Under a Bayesian framework with a normal prior on $\bm{\beta}$, this leads to straightfoward Gibbs samplers. For micobin regression, the same augmentation strategy can be adopted by conditioning on the mixing variable $\bm{\lambda}=(\lambda_{1},\dots,\lambda_{n})$.
Algorithms 1 and 2 describe blocked Gibbs samplers for cobin and micobin regression, where $X\in\mathbb{R}^{n\times p}$ is the design matrix. We assume a normal prior $\bm{\beta}\sim N_{p}(0,\Sigma_{\beta})$, some proper prior $\lambda\sim p_{\lambda}$ for cobin regression, and a beta prior $\psi\sim\mathrm{Beta}(a_{\psi},b_{\psi})$ for micobin regression. We set a large upper bound $L$ on $\lambda$, as the posterior probability of very large $\lambda$ is negligible in practice. The update steps for $\lambda$ and $\bm{\kappa}$ are blocked together, which improves mixing and avoids the evaluation of the density of KG in (11) when updating $\lambda$. The beta prior for $\psi$ in micobin regression leads to a conditionally conjugate update. In Appendix S.4, we describe the detailed derivations and the application of the Kolmogorov-Gamma augmentation to an EM algorithm for estimating the posterior mode, and discuss sampling strategies for varying dispersion micobin regression models (6) and spatial models (7).
Algorithm 1 One cycle of a blocked Gibbs sampler for cobin regression (2) with cobit link
1:Sample $\lambda$ from $\operatorname*{pr}(\lambda=l\mid\bm{\beta})\propto p_{\lambda}(l)\prod_{i=1}^{n}p_{\mathrm{cobin}}(y_{i};\bm{x}_{i}^{\mathrm{\scriptscriptstyle T}}\bm{\beta},l^{-1})$ among $\{1,\dots,L\}$
2:Sample $\kappa_{i}$ from $(\kappa_{i}\mid\lambda,\bm{\beta})\stackrel{{\scriptstyle\mathrm{ind}}}{{\sim}}\mathrm{KG}(\lambda,\bm{x}_{i}^{\mathrm{\scriptscriptstyle T}}\bm{\beta})$, $i=1,\dots,n$triangleright steps 1,2 jointly updates $(\lambda,\bm{\kappa})$
3:Sample $\bm{\beta}$ from $(\bm{\beta}\mid\lambda,\bm{\kappa})\sim N_{p}(\bm{m}_{\beta},V_{\beta})$, where
$V_{\beta}^{-1}=X^{\mathrm{\scriptscriptstyle T}}\hbox{diag}(\kappa_{1},\dots,\kappa_{n})X+\Sigma_{\beta}^{-1},\quad\bm{m}_{\beta}=V_{\beta}X^{\mathrm{\scriptscriptstyle T}}(y_{1}\lambda-0.5\lambda,\dots,y_{n}\lambda-0.5\lambda)^{\mathrm{\scriptscriptstyle T}}$
Algorithm 2 One cycle of a blocked Gibbs sampler for micobin regression (5) with cobit link
1:Sample $\lambda_{i}$ from $\operatorname*{pr}(\lambda_{i}=l\mid\bm{\beta},\psi)\propto l(1-\psi)^{l-1}p_{\mathrm{cobin}}(y_{i};\bm{x}_{i}^{\mathrm{\scriptscriptstyle T}}\bm{\beta},l^{-1})$ among $\{1,\dots,L\}$, $i=1,\dots,n$
2:Sample $\kappa_{i}$ from $(\kappa_{i}\mid\lambda_{i},\bm{\beta})\stackrel{{\scriptstyle\mathrm{ind}}}{{\sim}}\mathrm{KG}(\lambda_{i},\bm{x}_{i}^{\mathrm{\scriptscriptstyle T}}\bm{\beta})$, $i=1,\dots,n$triangleright steps 1,2 jointly updates $(\bm{\lambda},\bm{\kappa})$
3:Sample $\bm{\beta}$ from $(\bm{\beta}\mid\bm{\lambda},\bm{\kappa})\sim N_{p}(\bm{m}_{\beta},V_{\beta})$, where
$V_{\beta}^{-1}=X^{\mathrm{\scriptscriptstyle T}}\hbox{diag}(\kappa_{1},\dots,\kappa_{n})X+\Sigma_{\beta}^{-1},\quad\bm{m}_{\beta}=V_{\beta}X^{\mathrm{\scriptscriptstyle T}}(y_{1}\lambda_{1}-0.5\lambda_{1},\dots,y_{n}\lambda_{n}-0.5\lambda_{n})^{\mathrm{\scriptscriptstyle T}}$
4:Sample $\psi$ from $(\psi\mid\bm{\lambda})\sim\mathrm{Beta}(a_{\psi}+2n,b_{\psi}-n+\sum_{i=1}^{n}\lambda_{i})$triangleright steps 3,4 jointly updates $(\bm{\beta},\psi)$
The proposed Kolmogorov-Gamma augmentation scheme offers several advantages. Due to the conditionally normal likelihood, the algorithms can be trivially extended to mixed-effects models and more complex hierarchical models involving latent Gaussian structures. Moreover, by exploiting normal-normal conjugacy, the random and fixed effects can be updated jointly, which is especially important for spatial models (7) where the spatial random effect is often highly correlated with the intercept. In addition, a normal prior for $\bm{\beta}$ can be easily replaced by normal mixture priors, such as Laplace, Cauchy, or more broadly global-local shrinkage priors (Bhadra et al.,, 2016), simply by combining existing sampling methods developed for normal models.
Furthermore, we show that the proposed Gibbs sampler for cobin and micobin regression based on Kolmogorov-Gamma augmentation is uniformly ergodic, meaning that the Markov chain converges to the posterior geometrically fast in terms of total variation distance, uniformly at any initialization. This implies that the Markov chain mixes rapidly and guarantees the existence of central limit theorems for Monte Carlo averages of functions of $\bm{\beta}$.
Theorem 3.
The blocked Gibbs samplers presented in Algorithms 1 and 2 for cobin and micobin regressions are uniformly ergodic.
The proof follows a related approach to Choi and Hobert, (2013) based on a uniform minorization argument. Considering that popular approaches, such as Metropolis-adjusted Langevin or Hamiltonian Monte Carlo, rely on the assumption of log-concavity for their theoretical guarantees on fast mixing (Dwivedi et al.,, 2019; Chen et al.,, 2020), Theorem 3 is a strong result; for micobin regression the likelihood is not log-concave.
4 Sampling Kolmogorov-Gamma random variables
4.1 Kolmogorov-Gamma density
Developing a reliable and efficient sampling method for Kolmogorov-Gamma random variables is essential for our methodology. A naive approximation approach, based on truncating the sum of gamma random variables in Definition 3, is computationally inefficient and prone to truncation errors. In this section, we introduce an efficient method for exact sampling of $\mathrm{KG}(1,c)$ random variates based on the alternating series method of Devroye, (1986). One can then sample $\mathrm{KG}(b,c)$ variables by summing $b$ independent $\mathrm{KG}(1,c)$ variates.
We first describe the density of a $\mathrm{KG}(1,0)$ random variable, which can be easily derived from the two different density representations of the Kolmogorov distribution.
Proposition 2.
The $\mathrm{KG}(1,0)$ density has two different alternating series representations,
${p_{\textsc{kg}}}(x;1,0)=\sum_{n=0}^{\infty}(-1)^{n}a_{n}^{L}(x)$ $=\sum_{n=0}^{\infty}(-1)^{n}a^{R}_{n}(x),\quad x>0,$
$a_{n}^{L}(x)=\begin{cases}\frac{2}{\pi^{1/2}(2x)^{3/2}}\exp(-\frac{n^{2}}{8x})&(n\,\,\mathrm{odd})\\ \frac{(n+1)^{2}}{\pi^{1/2}(2x)^{5/2}}\exp(-\frac{(n+1)^{2}}{8x})&(n\,\,\mathrm{even})\end{cases},$ $\quad a_{n}^{R}(x)=4\pi^{2}(n+1)^{2}\exp\{-2\pi^{2}(n+1)^{2}x\}.$
Having two different density representations is crucial for developing a sampling scheme with the alternating series method. From Theorem 2, the density of $\mathrm{KG}(1,c)$ is ${p_{\textsc{kg}}}(x;1,c)=\left\{\sinh(c/2)/(c/2)\right\}\exp(-c^{2}x/2){p_{\textsc{kg}}}(x;1,0)$. Therefore, ${p_{\textsc{kg}}}(x;1,c)$ can also be represented as an alternating series ${p_{\textsc{kg}}}(x;1,c)=\sum_{n=0}^{\infty}(-1)^{n}a_{n}(x;c,t)$ where we define $a_{n}(x;c,t)$ as
$a_{n}(x;c,t)=\begin{cases}\left\{\sinh(c/2)/(c/2)\right\}\exp(-c^{2}x/2)a_{n}^{L}(x),&0<x<t,\\ \left\{\sinh(c/2)/(c/2)\right\}\exp(-c^{2}x/2)a_{n}^{R}(x),&t\leq x,\end{cases}$ (12)
for some suitable choice of cutoff point $t>0$, which will be discussed next.
4.2 Sampling KG(1,c) using alternating series method
The alternating series method (Devroye,, 1986) is an effective sampling algorithm when the target density $p(x)$ is computationally expensive to evaluate but can be approximated from above and below by a sequence of envelope functions $\{S_{m}(x)\}_{m=0}^{\infty}$, satisfying $S_{0}(x)>S_{2}(x)>\dots>p(x)>\dots>S_{3}(x)>S_{1}(x)$. With such an envelope function in hand, exact sampling of $X\sim p$ can be achieved with the following steps: (1) draw $X\sim q$ from the proposal distribution $q$, (2) generate $U\sim\mathrm{Unif}(0,Mq(X))$ where $\|p/q\|_{\infty}\leq M$, (3) repeat until $U\leq S_{m}(X)$ for odd $m$ or $U>S_{m}(X)$ for even $n$, (4) accept $X$ if $m$ is odd, repeat from (1) again if $m$ is even.
Similar to Devroye, (1986) and Polson et al., (2013), our choice of envelope function for sampling $\mathrm{KG}(1,c)$ is $S_{m}(x)=\sum_{n=0}^{m}(-1)^{n}a_{n}(x;c,t)$. For this choice to be a valid envelope function, $a_{n}(x;c,t)$ must be monotonically decreasing in $n$ for any fixed $x>0$ and parameter $c$. The following Lemma 1 shows the valid range of $t$, an intersection between the range of $x$ where $a_{n}^{L}(x)$ and $a_{n}^{R}(x)$ are both monotonically decreasing in $n$.
Lemma 1.
For any fixed $x>0$ and $c\in\mathbb{R}$, the sequence $a_{n}(x;c,t)$ defined in (12) is monotonically decreasing in $n$ when $\log(2)/(3\pi^{2})<t<0.25$, where $\log(2)/(3\pi^{2})\approx 0.0234$.
In what follows, we assume that $t$ satisfies Lemma 1; the optimal choice of $t$ will be discussed soon. Since $S_{0}(x)=a_{0}(x;c,t)>{p_{\textsc{kg}}}(x;1,c)$ for all $x$, a natural choice of proposal $q$ that ensures $\|{p_{\textsc{kg}}}(\cdot;1,c)/q(\cdot)\|_{\infty}\leq M$ for some $M$ is $q(x;c,t)=M^{-1}a_{0}(x;c,t)$. Plugging in $a_{0}^{L}(x)$ and $a_{0}^{R}(x)$ in (12), the proposal distribution with density $q(x;c,t)$ is
$X\sim\begin{cases}\mathrm{GIG}(-1.5,c^{2},1/4)1(0<X<t)\quad&\text{with prob. }A^{L}(c,t)/\{A^{L}(c,t)+A^{R}(c,t)\}\\ \mathrm{Exp}(c^{2}/2+2\pi^{2})1(t\leq X)&\text{with prob. }A^{R}(c,t)/\{A^{L}(c,t)+A^{R}(c,t)\}\\ \end{cases}$ (13)
where $\mathrm{GIG}(p,a,b)1(0<X\leq t)$ is a generalized inverse Gaussian (GIG) distribution with density proportional to $x^{p-1}\exp\{-(ax+b/x)/2\}$ truncated to $(0,t)$, $\mathrm{Exp}(a)1(t\leq X)$ is an exponential distribution with rate $a$ truncated to $[t,\infty)$, $A^{L}(c,t)=\int_{0}^{t}a_{0}(x;c,t)\mathrm{d}x$, and $A^{R}(c,t)=\int_{t}^{\infty}a_{0}(x;c,t)\mathrm{d}x$. Compared to the Pólya-Gamma case (Polson et al.,, 2013), which involves inverse Gaussian and exponential distributions in its proposal, our proposal is slightly more complicated, involving GIG and exponential distributions. Peña and Jauch, (2025) obtain results on the exact evaluation of the cdf and the sampling of GIG random variables that have a half-integer parameter $p=-1.5$. We use these results in calculating $A^{L}(c,t)$ and sampling truncated GIG random variables.
Our $\mathrm{KG}(1,c)$ sampling algorithm is highly efficient. We show this by first investigating how often a proposal $X\sim q$ from (13) is accepted, based on the expected number of outer loop iterations $M=\int_{0}^{\infty}a_{0}(x;c,t)\mathrm{d}x=A^{L}(c,t)+A^{R}(c,t)$, and then by illustrating that with very high probability, only few series terms $S_{m}(x)$ need to be computed in the inner loop in order to decide whether to accept or reject $X$.
Proposition 3.
The following statements hold:
  1. 1.
    The best cutoff point $t^{*}$ minimizing the expected number of outer loop iterations $M=A^{L}(c,t)+A^{R}(c,t)$ is independent of $c$; this value is $t^{*}\approx 0.050239$.
  2. 2.
    Using the best cutoff point $t^{*}$, $M$ is bounded above by $1.1456$ and the expected number of inner loop iterations is bounded above by $1.1275$ for any given $c$.
Compared to the Pólya-Gamma sampler implemented in the BayesLogit R package (Polson et al.,, 2013), which took $\sim 0.3$ seconds to generate 1 million samples from Pólya-Gamma(1,2), an Rcpp implementation of our Kolmogorov-Gamma sampler took $\sim 0.5$ seconds to generate 1 million samples from $\mathrm{KG}(1,2)$ in an Apple M1 CPU environment. Algorithm S.1 in Appendix S.2 describes the pseudocode and further details.
5 Simulation studies
5.1 Robustness of cobin regression estimator
We conducted a simulation study to empirically validate Proposition 1, focusing on the consistency and robustness of the cobin and beta regression estimators. We considered four different scenarios for the response distribution: (A) beta, (B) cobin, (C) mixture of beta and uniform distributions, called the beta rectangular distribution (Bayes et al.,, 2012), and (D) mixture of three beta distributions. Specifically, the densities of (C) and (D) with mean $\mu$ correspond to
$p_{\mathrm{brec}}(y;\mu,\alpha,\phi)$ $=w(\mu,\alpha)p_{\mathrm{beta}}\left(y;\frac{\mu-0.5+0.5w(\mu,\alpha)}{w(\mu,\alpha)},\phi\right)+1-w(\mu,\alpha),$ (14)
$p_{\mathrm{bmix}}(y;\mu,\phi)$ $=0.25p_{\mathrm{beta}}(y;\mu-\epsilon(\mu),\phi)+0.5p_{\mathrm{beta}}(y;\mu,\phi)+0.25p_{\mathrm{beta}}(y;\mu+\epsilon(\mu),\phi),$ (15)
for $y\in(0,1)$, where $p_{\mathrm{beta}}(y;\mu,\phi)$ corresponds to the beta density parametrized by mean $\mu$ and precision $\phi$ (sum of two beta shape parameters), $w(\mu,\alpha)=1-\alpha(1-|2\mu-1|)$ for some $\alpha\in(0,1)$, and $\epsilon(\mu)=\min(\mu,1-\mu)/2$. We also consider two different link functions (cobit, logit) with sample sizes $n\in\{100,400,1600\}$, resulting in $4\times 2\times 3=24$ different data generating scenarios. We consider two covariates including the intercept and set the true coefficients as $\bm{\beta}=(\beta_{0},\beta_{1})=(0,1)$. The non-intercept covariate is generated as $x_{i}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}N(0,\sigma_{x}^{2})$, where we set $\sigma_{x}^{2}=9$ for cobit and $\sigma_{x}^{2}=1$ for logit to account for the difference in scales between the link functions reflecting that $g^{\prime}_{\mathrm{cobit}}(0.5)$ and $g^{\prime}_{\mathrm{logit}}(0.5)$ differ by a factor of 3. For parameters not related to the mean, we set $\phi=8$ for beta, $\lambda=3$ for cobin, $(\alpha,\phi)=(0.2,10)$ for beta rectangular and $\phi=40$ for mixture of beta. For the beta rectangular, the choice of $\alpha=0.2$ ensures that the weight $w(\mu,\alpha)$ in (14) assigned to the beta distribution is greater than 0.8. We simulate 1000 replicated datasets for each data generation scenario. For each data set, we fit the cobin and beta regression models to find $\hat{\bm{\beta}}$ with unknown dispersion parameters, using the correct link function but not necessarily the true data generating distribution. We used iteratively reweighted least squares to find $\hat{\bm{\beta}}$ for cobin regression based on the stats package in R (R Core Team,, 2024), and we utilized the betareg R package (Cribari-Neto and Zeileis,, 2010) for the beta regression estimate.
Table 1: Analysis of beta and cobin regression estimates, better results are highlighted in bold. Results are based on 1000 replicates under the correct mean structure but possibly misspecified response distributions.
Beta Cobin Beta rectangular Mixture of beta
Link Method $n$ Bias RMSE Bias RMSE Bias RMSE Bias RMSE
cobit beta regression $100$ 0.004 0.106 -0.024 0.099 -0.061 0.153 -0.030 0.094
$400$0.003 0.054 -0.030 0.057 -0.069 0.099 -0.037 0.058
$1600$ 0.002 0.026 -0.030 0.038 -0.076 0.084 -0.036 0.042
cobin regression $100$ 0.005 0.113 0.003 0.099 0.013 0.134 0.006 0.092
$400$0.002 0.056 -0.001 0.049 0.006 0.070 -0.001 0.046
$1600$ 0.002 0.027 0.000 0.023 -0.001 0.035 0.001 0.022
logit beta regression $100$ 0.003 0.084 -0.043 0.080 -0.054 0.117 -0.041 0.074
$400$ 0.000 0.042 -0.047 0.059 -0.059 0.080 -0.046 0.055
$1600$ 0.000 0.021 -0.045 0.048 -0.062 0.068 -0.046 0.048
cobin regression $100$ 0.015 0.101 0.005 0.066 0.020 0.116 0.005 0.067
$400$ 0.004 0.051 0.000 0.035 0.007 0.062 0.000 0.033
$1600$ 0.000 0.026 0.001 0.016 0.001 0.032 0.001 0.016
The result is summarized in Table 1 in terms of bias and root mean square error (RMSE) of $\hat{\beta}_{1}$. As $n$ increases, the estimates from the cobin regression model exhibit a decreasing bias for any data-generating scenario, supporting the consistency described in Proposition 1. This stands in contrast to the results from the beta regression model, which show persistent bias and inconsistency under data-generating scenarios other than beta. For cobin regression results, the RMSE decreases proportionally to $n^{-1/2}$ for any data-generating scenario, which aligns with the asymptotic normality of $\hat{\bm{\beta}}$ in correctly specified or misspecified cases (Gourieroux et al.,, 1984).
5.2 Robustness and scalability under spatial regression models
Next, we consider a spatial regression simulation study to analyze the robustness of cobin and micobin regression models against model misspecification, and highlight the computational benefits of the Kolmogorov-Gamma augmentation, particularly for latent Gaussian models. For data generation, we choose spatial locations uniformly at random from $[0,1]^{2}$ as training and test locations with sizes $(n_{\mathrm{train}},n_{\mathrm{test}})\in\{(200,50),(400,100)\}$. Then, spatial random effects are generated from a mean zero Gaussian process (GP) with exponential kernel $\operatorname*{cov}\{u(s),u(s^{\prime})\}=\sigma_{u}^{2}\exp(-\|s-s^{\prime}\|_{2}/\rho)$, where we set $\sigma_{u}^{2}=1$ and $\rho\in\{0.1,0.2\}$. For the fixed effect terms, we consider two covariates including the intercept, generate a non-intercept covariate as $x(s_{i})\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}N(0,3^{2})$, and set true coefficients as $\bm{\beta}=(\beta_{0},\beta_{1})=(0,1)$. The responses $y(s_{i})$ are generated from a beta rectangular distribution (14) with cobit link $\mu(s_{i})=g_{\mathrm{cobit}}^{-1}\{x(s_{i})\beta_{1}+u(s_{i})\}$, $\alpha=0.2$, and $\phi=10$. The beta rectangular response distribution leads to data that are occasionally sampled from the uniform distribution, reflecting potential outliers often encountered in realistic scenarios. This data generation process is repeated 100 times.
Based on the training set with size $n_{\mathrm{train}}$, we fit three different spatial regression models (beta, cobin, micobin) with correct mean structure (7) but none of the response distributions are correctly specified. We set a normal prior on the regression coefficients $\bm{\beta}\sim N_{2}(\bm{0}_{2},100^{2}\mathbf{I}_{2})$ and a half-Cauchy prior on the spatial random effects standard deviation $\sigma_{u}$. The spatial range parameter $\rho\in\{0.1,0.2\}$ is fixed at the true value. For spatial cobin and micobin regression models, we employ a blocked Gibbs sampler with Kolmogorov-Gamma augmentation. For the spatial beta regression model, we use Stan to carry out posterior inference using the No-U-Turn Sampler algorithm (Carpenter et al.,, 2017). We run a total of 6,000 MCMC iterations and record wall-clock running time, with the first 1,000 samples discarded as burn-in. For further details of the simulation settings, including priors for dispersion parameters, see Appendix S.5.
Table 2: Spatial regression simulation results under misspecified response distribution based on 100 replicates. Better results are highlighted in bold except for sampling performance.
Inference $(\hat{\beta}_{1})$ Prediction Sampling $(\bm{\beta})$
$\rho$ Method $(n_{\mathrm{train}},n_{\mathrm{test}})$ Bias RMSE negtestLL MSPE×$\times 10^{2}$ mESS time (min)
beta regression $(200,50)$0.048 0.118 -0.325 0.427 919.8 44.5
$(400,100)$0.052 0.089 -0.354 0.345 978.7 437.7
0.1 cobin regression $(200,50)$ 0.005 0.093 -0.340 0.388 2791.3 2.0
$(400,100)$ 0.005 0.067 -0.372 0.323 3220.9 11.2
micobin regression $(200,50)$ 0.034 0.099 -0.367 0.373 1908.4 2.4
$(400,100)$ 0.037 0.074 -0.394 0.312 2137.5 11.7
beta regression $(200,50)$0.065 0.120 -0.320 0.329 1187.2 96.3
$(400,100)$0.052 0.095 -0.350 0.248 808.0 933.4
0.2 cobin regression $(200,50)$ 0.000 0.088 -0.346 0.306 3366.0 2.2
$(400,100)$ 0.013 0.078 -0.370 0.233 3663.9 12.1
micobin regression $(200,50)$ 0.039 0.092 -0.373 0.293 2265.3 2.2
$(400,100)$ 0.050 0.091 -0.395 0.226 2575.4 12.7
Monte Carlo standard errors are all less than 0.015 for negtestLL, 0.013 for MSPE, 127.2 for mESS.
Table 2 summarizes the results. First, with respect to the posterior mean estimate $\hat{\beta}_{1}$ of the fixed effect coefficient, cobin regression consistently produces the lowest bias and RMSE in all scenarios, in accordance with the previous simulation results. In contrast, beta and micobin regression estimates exhibit bias, with the magnitude of these biases increasing as spatial dependence becomes stronger. Second, to evaluate predictive performance under model misspecification, we compare the negative test log-likelihood conditional on random effects (negtestLL) and mean square prediction error (MSPE) based on test data. Micobin regression outperforms the others, while beta regression performs the worst on both metrics, highlighting the robustness of micobin regression in prediction. Finally, in terms of computational efficiency, cobin and micobin achieve a significantly higher multivariate effective sample size (mESS) (Vats et al.,, 2019) of $\bm{\beta}$ per unit time compared to beta regression. These findings demonstrate the scalability and robustness of cobin and micobin regression in estimation and prediction, respectively.
6 Benthic macroinvertebrate multimetric index of U.S. lakes
As an illustrative application of cobin and micobin regression models with random effects, we analyze the benthic macroinvertebrate multimetric index (MMI) of US lakes and the association with lake watershed covariates. MMI, also known as an index of biotic integrity, is a standard quantitative measure for the bioassessment of macroinvertebrate assemblages (Karr,, 1991; Stoddard et al.,, 2008) that integrates various attributes of the assemblage (e.g. taxonomic composition and richness). Higher MMI values indicate a healthier and more diverse benthic macroinvertebrate community. We consider MMI data from the 2017 National Lake Assessment Survey (NLA), which covers about 1,000 lakes in the conterminous US; see the left panel of Figure 1. We refer to the 2017 NLA survey technical report for details on MMI (U.S. Environmental Protection Agency,, 2022).
We are interested in understanding how the biotic integrity of lake ecosystems measured by MMI is associated with natural and human-related lake watershed characteristics, as well as in predicting the MMI of unsurveyed lakes. We consider LakeCat data (Hill et al.,, 2018) covering more than 380,000 US lakes, containing various natural and anthropogenic watershed covariates. For illustrative purposes, we select 7 watershed covariates that are highly important in the analysis of lake eutrophication (Hill et al.,, 2018, Fig. 7), as well as 2 additional covariates (manure application and urban land cover); see Table S.3 in Appendix S.6 for description. Since all covariates were heavily right-skewed and exhibited limited variation around the mean, we applied the transformation $x\mapsto\log_{2}(x+1)$ to reduce skewness and mitigate the influence of very large values.
Among the 2017 NLA survey data, a total of 950 lakes remained after removing those missing MMI and/or covariates. One lake had an MMI value of 0, and we removed it to allow a comparison with spatial beta and cobin regression models; the spatial micobin regression result with original data is available in Appendix S.6.
For predictive assessments, we selected 55,215 lakes from the LakeCat dataset, focusing on those with surface areas greater than 40,000 m2. In the spatial model with cobit link (7), making probabilistic predictions for these 55,215 lakes using a traditional GP prior is computationally prohibitive. Thus, we employ a nearest neighbor Gaussian process (NNGP) (Datta et al.,, 2016) on the spatial random effect $u(\cdot)$. We use a prior and algorithm similar to that in Section 5.2 for spatial cobin and micobin regression, and Stan for spatial beta regression; see Appendix S.6 for details. We ran three chains for each model for a total of 6,000 iterations per chain, discarding the first 1,000 samples from each as burn-in. The convergence diagnosis did not indicate any problems.
Table 3: Comparison of MMI data analysis results under three different models. Variables whose 95% credible intervals do not include zero are highlighted in bold. Asterisks indicate variables for which the 95% credible interval does not contain zero after removing two influential observations (See Figure 4).
($n=949$) Beta regression Cobin regression Micobin regression
Variable Estimate 95% CI Estimate 95% CI Estimate 95% CI
Intercept -2.363 (-4.160, -0.553) -2.106 (-3.859, -0.345) -1.797 (-3.551, -0.085)
agkffact -2.586 (-5.584, 0.330) -2.888 (-5.714, -0.003) -3.457 (-6.113, -0.800)
bfi 0.343 (0.016, 0.672) 0.293 (-0.022, 0.614) 0.229 (-0.082, 0.548)
cbnf 0.165 (-0.081, 0.412) 0.182 (-0.055, 0.420) 0.191 (-0.035, 0.425)
conif 0.081 (-0.002, 0.164) 0.093 (0.011, 0.176) 0.123 (0.044, 0.203)
crophay -0.079 (-0.250, 0.091) -0.063 (-0.231, 0.106) -0.054 (-0.213, 0.105)
fert -0.073 (-0.310, 0.158) -0.092 (-0.323, 0.132) -0.082 (-0.300, 0.138)
manure -0.048 (-0.202, 0.102) -0.036 (-0.182, 0.115) -0.029 (-0.173, 0.118)
pestic97 -0.014 (-0.106, 0.075) -0.021 (-0.108, 0.067) -0.025 (-0.108, 0.059)
urbmdhi -0.181 (-0.288, -0.076) -0.170 (-0.273, -0.067) -0.142 (-0.243, -0.041)
agkffact, soil erodibility factor; bfi, base flow index; cbnf, cultivated biological N fixation; conif, coniferous forest cover; crophay, crop/hay land cover; fert, synthetic N fertilizer use; manure, manure application; pestic97, 1997 pesticide use; urbmdhi, medium/high-density urban land cover. All variables are $\log_{2}(x+1)$ transformed. See Table S.3 in Appendix S.6 for a detailed description.
Figure 3: Predicted MMI at 55,215 lakes from the spatial micobin regression model in terms of $E\{Y(s^{*})\mid X(s^{*})\}$ at unsampled location $s^{*}$. (Left) Posterior predictive mean of $E\{Y(s^{*})\mid X(s^{*})\}$. (Right) Posterior predictive standard deviation of $E\{Y(s^{*})\mid X(s^{*})\}$.
The estimated fixed-effect coefficients are summarized in Table 3. Our results show that decreased biotic integrity (low MMI) in US lakes is associated with high soil erodibility, low base flow index, low coniferous forest cover, and high urban land cover in the lake watershed area. The signs are generally sensible, and the variables selected based on 95% credible intervals coincide between cobin and micobin, but differ with beta regression. In terms of computation, both spatial cobin and micobin regressions outperform spatial beta regression in terms of mESS of $\bm{\beta}$ per unit time by more than a factor of 20. Regarding WAIC (Gelman et al.,, 2014) conditional on random effects, the beta, cobin, and micobin regression models yield -1093.4, -1103.5, and -1119.3, respectively, suggesting that micobin regression achieves the best predictive accuracy among the three models. Figure 3 illustrates the predicted MMI at 55,215 lakes, with the right panel reflecting increased uncertainty in southern California and western Texas due to sparse MMI data.
We further analyze quantile residuals (Dunn and Smyth,, 1996) to assess goodness-of-fit, with the model parameters fixed at the posterior mean (posterior median of $\lambda$ for cobin). Figure 4 shows that beta and cobin regression exhibit a lack of fit in the left tail, while micobin regression captures the left tail accurately but exhibits a slight under-estimation of the right tail. Further investigation reveals that the two influential observations with the lowest quantile residuals are identical for all three models, with MMI values of 0.02 and 0.021. We repeated the analysis, removing these observations. After removal, the variables selected based on 95% credible intervals that did not overlap zero changed for beta regression, while the results for the cobin and micobin regression remained the same. This is consistent with the beta regression score function being unbounded in $y$, leading to brittle results. See Appendix S.6 for details. Overall, the results suggest that the cobin and micobin regressions are more robust to observations near the boundaries, exhibit significantly better scalability, and achieve better predictive performance.
Figure 4: Comparison of quantile residuals for goodness-of-fit assessment, along with two observations corresponding to the lowest quantiles. The red line corresponds to the $y=x$ line.
7 Discussion
Beyond generalized linear (mixed) models, we anticipate that cobin and micobin distributions can be naturally incorporated into a more diverse family of models with continuous proportional data, such as tree ensembles (Schmid et al.,, 2013) or deep generative models (Loaiza-Ganem and Cunningham,, 2019). Developments of scalable inference methods are key to enabling such extensions with massive amounts of data. Similar to how Pólya-Gamma augmentation is connected with variational inference for logistic models (Durante and Rigon,, 2019), we hope that the proposed Kolmogorov-Gamma augmentation provides insight for the future development of approximate Bayesian inference methods, in addition to facilitating inference using MCMC.
Acknowledgement
This research was partially supported by the National Institutes of Health (grant ID R01ES035625), by the European Research Council under the European Union’s Horizon 2020 research and innovation programme (grant agreement No 856506), by the National Science Foundation (NSF IIS-2426762), and by the Office of Naval Research (N00014-21-1-2510).
Software
Code to reproduce the analyses is available at https://github.com/changwoo-lee/cobin-reproduce.
Appendices
Section S.1 contains proofs of statements presented in the main article. Section S.2 contains pseudocode of the Kolomogorov-Gamma $(1,c)$ distribution and additional sampling details. Section S.3 contains a detailed derivation of cobin as an exponential dispersion model, density and cumulative distribution functions of cobin and micobin, and a discussion on the link and variance functions. Section S.5 contains derivations of Gibbs samplers as well as an EM algorithm. Finally, Section S.6 gives further information about benthic macroinvertebrate multimetric index data, analysis settings, and additional results.
S.1 Proofs
S.1.1 Proof of Theorem 2
Proof.
The case when $c=0$ is trivial, thus assume $c\neq 0$. From the proof of Theorem 1, we know that $E(e^{-c^{2}\kappa/2})=\left\{\sinh(c/2)/(c/2)\right\}^{-b}$ for $\kappa\sim\mathrm{KG}(b,0)$, which corresponds to the normalizing constant of an exponential tilted density. Thus, the exponential tilted distribution has density
$\left\{\sinh(c/2)/(c/2)\right\}^{b}\exp(-c^{2}\kappa/2){p_{\textsc{kg}}}(\kappa;b,0).$ (S.1)
To show that the Laplace transformation of (S.1) coincides with the Laplace transformation of a $\mathrm{KG}(b,c)$ random variable defined as the infinite convolution of gamma distributions, let
$\int_{0}^{\infty}e^{-t\kappa}\frac{\{\sinh(c/2)\}^{b}}{(c/2)^{b}}e^{-c^{2}\kappa/2}{p_{\textsc{kg}}}(\kappa;b,0)\mathrm{d}\kappa$ =$=\frac{\{\sinh(c/2)\}^{b}}{(c/2)^{b}}\int e^{-(-0.5c^{2}+t)\kappa}{p_{\textsc{kg}}}(\kappa;b,0)d\kappa$
$=\frac{[\sinh\{(c^{2}/4)^{1/2}\}]^{b}}{\{(c^{2}/4)^{1/2}\}^{b}}\frac{[\{(0.5c^{2}+t)/2\}^{1/2}]^{b}}{(\sinh{[\{(0.5c^{2}+t)/2\}^{1/2}}])^{b}}$
$=\prod_{k=1}^{\infty}\left\{\frac{\frac{2k^{2}\pi^{2}+c^{2}/2}{2k^{2}\pi^{2}}}{\frac{2k^{2}\pi^{2}+(0.5c^{2}+t)}{2k^{2}\pi^{2}}}\right\}^{b}$
$=\prod_{k=1}^{\infty}(1+d_{k}^{-1}t)^{-b}$
where $d_{k}=2k^{2}\pi^{2}+c^{2}/2$. This corresponds to the Laplace transformation of the infinite sum of independent $\mathrm{Gamma}(b,1)$ distributions scaled by $d_{k}^{-1}$ for $k=1,\dots$, which completes the proof. ∎
S.1.2 Proof of Theorem 3
First, define $\mathrm{sinhc}(x)\coloneqq\sinh(x)/x$ for $x\neq 0$ and $\mathrm{sinhc}(0)\coloneqq 1$. We first present two technical lemmas, where Lemma S.1 is the same as Lemma 3.1 of Choi and Hobert, (2013).
Lemma S.1.
If $A$ is a symmetric nonnegative definite matrix, all eigenvalues of $(I+A)^{-1}$ are in $(0,1]$ and thus $\mathbf{z}^{\mathrm{\scriptscriptstyle T}}(I+A)^{-1}\mathbf{z}\leq\mathbf{z}^{\mathrm{\scriptscriptstyle T}}\mathbf{z}$ for any vector $\mathbf{z}$. Also, $I-(I+A)^{-1}$ is nonnegative definite.
Lemma S.2.
For $a,b\geq 0$, $\mathrm{sinhc}(a+b)\leq 2\mathrm{sinhc}(a)\cosh(b)$.
Proof.
It trivially holds when $a=0$ or $b=0$. When $a,b>0$, by expanding $\sinh(a+b)=\sinh(a)\cosh(b)+\cosh(a)\sinh(b)$ and multiplying both side by $a+b$, it is equivalent to showing the inequality $\coth(a)\tanh(b)\leq 1+2b/a$ for $a,b>0$. In other words, it suffices to show that for any given $a>0$, $f(x)=\coth(a)\tanh(x)-1-2x/a\leq 0$ for all $x>0$. Consider two linear functions $g_{1}(x)=x-1$ and $g_{2}(x)=\coth(a)-1-2x/a$, with $g_{1}$ increasing and $g_{2}$ decreasing. It can be easily checked that $g_{1}(x)\geq f(x)$ and $g_{2}(x)\geq f(x)$ for any $x>0$, thus $\min(g_{1}(x),g_{2}(x))\geq f(x)$ for any $x>0$. The proof is completed by confirming $g_{1}(x)$ and $g_{2}(x)$ intersects at $x^{*}=a\coth(a)/(a+2)$ with $g_{1}(x^{*})=g_{2}(x^{*})=a\coth(a)/(a+2)-1\leq 0$ for any $a>0$. ∎
Now we provide the formal statement of the theorem. We denote $P^{t}$ as a $t$-step transition kernel, $\|\nu_{1}-\nu_{2}\|_{\mathrm{TV}}$ as a total variation distance between probability measures $\nu_{1}$ and $\nu_{2}$. Let $\bm{\Theta}$ be a set of parameters of the model and $\Pi(\cdot)$ be a posterior of $\bm{\Theta}$. We say the Markov chain $\{\bm{\Theta}^{(m)}\}_{m=0}^{\infty}$ is uniformly ergodic if there exist a constant $M>0$ and $\rho\in[0,1)$, both independent of initial state $\bm{\Theta}^{(0)}$, such that $\|P^{t}(\bm{\Theta}^{(0)},\cdot)-\Pi(\cdot)\|_{\mathrm{TV}}\leq M\rho^{t}$ for all $t\geq 1$.
When parameters are partitioned into two blocks $\bm{\Theta}=\bm{\Theta}_{1}\cup\bm{\Theta}_{2}$, $\bm{\Theta}_{1}\cap\bm{\Theta}_{2}=\emptyset$ and a Gibbs sampler iteratively updates between $p(\bm{\Theta}_{1}\mid\bm{\Theta}_{2})$ and $p(\bm{\Theta}_{2}\mid\bm{\Theta}_{1})$, showing the uniform ergodicity of either $\{\bm{\Theta}_{1}^{(m)}\}_{m=0}^{\infty}$ or $\{\bm{\Theta}_{2}^{(m)}\}_{m=0}^{\infty}$ is sufficient for the uniform ergodicity of $\{\bm{\Theta}^{(m)}\}_{m=0}^{\infty}$; see Roberts and Rosenthal, (2001) and also Bhattacharya et al., (2021). The conditioning on data $\bm{y}$ is always assumed and suppressed from the notation for simplicity. Recall that Algorithm 1 for cobin regression consists of two blocks $\bm{\Theta}_{1}=(\lambda,\bm{\kappa})$ and $\bm{\Theta}_{2}=\bm{\beta}$. Also, Algorithm 2 for micobin regression consists of two blocks $\bm{\Theta}_{1}=(\bm{\lambda},\bm{\kappa})$ and $\bm{\Theta}_{2}=(\bm{\beta},\psi)$. To end with, we establish uniform ergodicity by showing that the marginal chain $\{\bm{\Theta}_{2}^{(m)}\}_{m=0}^{\infty}$ is uniformly ergodic for both Algorithms, under a mean zero normal prior for coefficient $\bm{\beta}\sim N_{p}(\bm{0},\Sigma_{\beta})$ (and beta prior $\psi\sim\mathrm{Beta}(a_{\psi},b_{\psi})$ for micobin) and with some large upper bound $L$ of $\lambda$.
The proof strategy is based on the establishment of a uniform minorization condition, also known as a Doeblin condition (Rosenthal,, 1995; Jones and Hobert,, 2001). Our approach is structurally similar to Choi and Hobert, (2013), but now involves Kolmogorov-Gamma variables instead of Polya-Gamma, and additionally involves $\bm{\lambda}=(\lambda_{1},\dots,\lambda_{n})$ as well as $\psi$ for micobin. Letting $k(\bm{\Theta}_{2}\mid\bm{\Theta}_{2}^{*})$ be a Markov transition density, it is sufficient to show that there is a constant $\delta^{\star}>0$ and a probability density function $q(\bm{\Theta}_{2})$, which does not depend on $\bm{\Theta}_{2}^{*}$, such that $k(\bm{\Theta}_{2}\mid\bm{\Theta}_{2}^{*})\geq\delta^{\star}q(\bm{\Theta}_{2})$ for any $\bm{\Theta}_{2},\bm{\Theta}_{2}^{*}$.
We first focus on Algorithm 2 for micobin regression, where $\bm{\Theta}_{2}=(\bm{\beta},\psi)$, and then specialize to cobin afterwards. Let $\bm{\Theta}_{2}^{*}=(\bm{\beta}^{*},\psi^{*})$ be a parameter of the previous step. The Markov transition density of $\bm{\Theta}_{2}$ (marginalizing out $\bm{\Theta}_{1}$) is
$k(\bm{\Theta}_{2}\mid\bm{\Theta}^{*}_{2})=\sum_{\bm{\lambda}\in\{1,\dots,L\}^{n}}p(\psi\mid\bm{\lambda})\left\{\int_{(0,\infty)^{n}}p(\bm{\beta}\mid\bm{\kappa},\bm{\lambda})p(\bm{\kappa}\mid\bm{\beta}^{*},\bm{\lambda})\mathrm{d}\bm{\kappa}\right\}p(\bm{\lambda}\mid\bm{\beta}^{*},\psi^{*})$ (S.2)
We aim to establish a lower bound of the curly bracket term in (S.2) that is independent of $\bm{\beta}^{*}$. This is achieved through two steps.
Proposition S.1.
Denoting $\mathbf{s}=\mathbf{s}(\bm{\lambda})=\Sigma_{\beta}^{1/2}X^{\mathrm{\scriptscriptstyle T}}\tilde{\bm{y}}$ where $\tilde{\bm{y}}\in\mathbb{R}^{n}$ is a vector with $i$th element $(y_{i}\lambda_{i}-\lambda_{i}/2)$,
$p(\bm{\beta}\mid\bm{\kappa},\bm{\lambda})\geq\frac{1}{(2\pi)^{p/2}|\Sigma_{\beta}|^{1/2}}\exp\left(-\frac{1}{2}\bm{\beta}^{\mathrm{\scriptscriptstyle T}}\Sigma_{\beta}^{-1}\bm{\beta}-\frac{1}{2}\mathbf{s}^{\mathrm{\scriptscriptstyle T}}\mathbf{s}+\mathbf{s}^{\mathrm{\scriptscriptstyle T}}\Sigma_{\beta}^{-1/2}\bm{\beta}\right)\prod_{i=1}^{n}\exp\left\{-\frac{(\bm{x}_{i}^{\mathrm{\scriptscriptstyle T}}\bm{\beta})^{2}}{2}\kappa_{i}\right\}$
Proof.
Letting $K=\hbox{diag}(\bm{\kappa})$ and $V_{\beta}=(X^{\mathrm{\scriptscriptstyle T}}KX+\Sigma_{\beta}^{-1})^{-1}$, we have $|V_{\beta}|\leq|\Sigma_{\beta}|$ by Lemma S.1 since $|V_{\beta}|=|\Sigma_{\beta}||(\tilde{X}^{\mathrm{\scriptscriptstyle T}}K\tilde{X}+I_{p})^{-1}|$ where $\tilde{X}=X\Sigma_{\beta}^{1/2}$. Next, for $\mathbf{m}=V_{\beta}X^{\mathrm{\scriptscriptstyle T}}\tilde{\bm{y}}$ we have $\mathbf{m}^{\mathrm{\scriptscriptstyle T}}V_{\beta}^{-1}\mathbf{m}\leq\mathbf{s}^{\mathrm{\scriptscriptstyle T}}\mathbf{s}$, which follows from $\mathbf{m}^{\mathrm{\scriptscriptstyle T}}V_{\beta}^{-1}\mathbf{m}=(\tilde{X}\tilde{\bm{y}})^{\mathrm{\scriptscriptstyle T}}(\tilde{X}^{\mathrm{\scriptscriptstyle T}}K\tilde{X}+I_{p})^{-1}(\tilde{X}\tilde{\bm{y}})\leq\mathbf{s}^{\mathrm{\scriptscriptstyle T}}\mathbf{s}$ from Lemma S.1. Using two inequalities and $\mathbf{m}^{\mathrm{\scriptscriptstyle T}}V_{\beta}^{-1}=\mathbf{s}^{\mathrm{\scriptscriptstyle T}}\Sigma_{\beta}^{-1/2}$, we have
$p(\bm{\beta}\mid\bm{\kappa},\bm{\lambda})$ $=(2\pi)^{-p/2}|V_{\beta}|^{-1/2}\exp(-(\bm{\beta}-\mathbf{m})^{\mathrm{\scriptscriptstyle T}}V_{\beta}^{-1}(\bm{\beta}-\mathbf{m})/2)$
$\geq(2\pi)^{-p/2}|\Sigma_{\beta}|^{-1/2}\exp\left(-\frac{1}{2}\bm{\beta}^{\mathrm{\scriptscriptstyle T}}\Sigma_{\beta}^{-1}\bm{\beta}-\frac{1}{2}\mathbf{s}^{\mathrm{\scriptscriptstyle T}}\mathbf{s}+\mathbf{s}^{\mathrm{\scriptscriptstyle T}}\Sigma_{\beta}^{-1/2}\bm{\beta}\right)\prod_{i=1}^{n}\exp\left\{-\frac{(\bm{x}_{i}^{\mathrm{\scriptscriptstyle T}}\bm{\beta})^{2}}{2}\kappa_{i}\right\}$
Proposition S.2.
We have a lower bound of the curly bracket term in (S.2):
$\int_{(0,\infty)^{n}}p(\bm{\beta}\mid\bm{\kappa},\bm{\lambda})p(\bm{\kappa}\mid\bm{\beta}^{*},\bm{\lambda})\mathrm{d}\bm{\kappa}\geq\delta(\bm{\lambda})N_{p}(\bm{\beta};\bm{\mu}^{\star}(\bm{\lambda}),\Sigma^{\star})$
where $\Sigma^{\star}=(\frac{1}{2}X^{\mathrm{\scriptscriptstyle T}}X+\Sigma_{\beta}^{-1})^{-1}$, $\bm{\mu}^{\star}=\bm{\mu}^{\star}(\bm{\lambda})=(\frac{1}{2}X^{\mathrm{\scriptscriptstyle T}}X+\Sigma_{\beta}^{-1})^{-1}\Sigma_{\beta}^{-1/2}\mathbf{s}(\bm{\lambda})$, and $\delta(\bm{\lambda})=2^{-\sum_{i}\lambda_{i}}e^{-\sum_{i=1}^{n}\lambda_{i}^{2}/4}|\Sigma^{\star}|^{1/2}|\Sigma_{\beta}|^{-1/2}\exp\left(-\frac{1}{2}\mathbf{s}^{\mathrm{\scriptscriptstyle T}}(I_{p}-(\Sigma_{\beta}^{1/2}X^{\mathrm{\scriptscriptstyle T}}X\Sigma_{\beta}^{1/2}/2+I_{p})^{-1})\mathbf{s}\right)$.
Proof.
First, we have $p(\bm{\kappa}\mid\bm{\beta}^{*},\bm{\lambda})=\prod_{i=1}^{n}\mathrm{sinhc}(\bm{x}_{i}^{\mathrm{\scriptscriptstyle T}}\bm{\beta}^{*}/2)^{\lambda_{i}}\exp(-(\bm{x}_{i}^{\mathrm{\scriptscriptstyle T}}\bm{\beta}^{*})^{2}\kappa_{i}/2){p_{\textsc{kg}}}(\kappa_{i};\lambda_{i},0)$. Then, with Proposition S.1, the integrand has a lower bound
$p(\bm{\beta}\mid\bm{\kappa},\bm{\lambda})p(\bm{\kappa}\mid\bm{\beta}^{*},\bm{\lambda})$ $\geq(2\pi)^{-p/2}|\Sigma_{\beta}|^{-1/2}\exp\left(-\frac{1}{2}\bm{\beta}^{\mathrm{\scriptscriptstyle T}}\Sigma_{\beta}^{-1}\bm{\beta}-\frac{1}{2}\mathbf{s}^{\mathrm{\scriptscriptstyle T}}\mathbf{s}+\mathbf{s}^{\mathrm{\scriptscriptstyle T}}\Sigma_{\beta}^{-1/2}\bm{\beta}\right)$
×$\times\prod_{i=1}^{n}\mathrm{sinhc}\left(\frac{\bm{x}_{i}^{\mathrm{\scriptscriptstyle T}}\bm{\beta}^{*}}{2}\right)^{\lambda_{i}}\exp\left(-\frac{\left\{(\bm{x}_{i}^{\mathrm{\scriptscriptstyle T}}\bm{\beta}^{*})^{2}+(\bm{x}_{i}^{\mathrm{\scriptscriptstyle T}}\bm{\beta})^{2}\right\}\kappa_{i}}{2}\right){p_{\textsc{kg}}}(\kappa_{i};\lambda_{i},0)$
Analyzing the terms involving $\kappa_{i}$, after integration,
$\int_{0}^{\infty}\exp\left(-\frac{\left\{(\bm{x}_{i}^{\mathrm{\scriptscriptstyle T}}\bm{\beta}^{*})^{2}+(\bm{x}_{i}^{\mathrm{\scriptscriptstyle T}}\bm{\beta})^{2}\right\}\kappa_{i}}{2}\right){p_{\textsc{kg}}}(\kappa_{i};\lambda_{i},0)\mathrm{d}\kappa_{i}$ =$=\left\{\mathrm{sinhc}\left(\frac{\sqrt{|\bm{x}_{i}^{\mathrm{\scriptscriptstyle T}}\bm{\beta}^{*}|^{2}+|\bm{x}_{i}^{\mathrm{\scriptscriptstyle T}}\bm{\beta}|^{2}}}{2}\right)\right\}^{-\lambda_{i}}$
$\geq\left\{\mathrm{sinhc}\left(\frac{|\bm{x}_{i}^{\mathrm{\scriptscriptstyle T}}\bm{\beta}^{*}|}{2}+\frac{|\bm{x}_{i}^{\mathrm{\scriptscriptstyle T}}\bm{\beta}|}{2}\right)\right\}^{-\lambda_{i}}$
$\geq\left\{2\mathrm{sinhc}\left(\frac{|\bm{x}_{i}^{\mathrm{\scriptscriptstyle T}}\bm{\beta}^{*}|}{2}\right)\cosh\left(\frac{|\bm{x}_{i}^{\mathrm{\scriptscriptstyle T}}\bm{\beta}|}{2}\right)\right\}^{-\lambda_{i}}$
where the first inequality is due to $\sqrt{a+b}\leq\sqrt{a}+\sqrt{b}$ for $a,b\geq 0$, combining with $\mathrm{sinhc}(x)^{-\lambda_{i}}$ is a decreasing function for $x\geq 0$, and the second inequality is by Lemma S.2. Thus, we have a lower bound of the integral that is independent of $\bm{\beta}^{*}$,
$\int_{(0,\infty)^{n}}\prod_{i=1}^{n}\mathrm{sinhc}\left(\frac{\bm{x}_{i}^{\mathrm{\scriptscriptstyle T}}\bm{\beta}^{*}}{2}\right)^{\lambda_{i}}\exp\left(-\frac{\left\{(\bm{x}_{i}^{\mathrm{\scriptscriptstyle T}}\bm{\beta}^{*})^{2}+(\bm{x}_{i}^{\mathrm{\scriptscriptstyle T}}\bm{\beta})^{2}\right\}\kappa_{i}}{2}\right){p_{\textsc{kg}}}(\kappa_{i};\lambda_{i},0)\mathrm{d}\bm{\kappa}$
$\geq$ $2^{-\sum_{i}\lambda_{i}}\left[\prod_{i=1}^{n}\cosh\left(\frac{|\bm{x}_{i}^{\mathrm{\scriptscriptstyle T}}\bm{\beta}|}{2}\right)^{-\lambda_{i}}\right]$
$\geq$ $2^{-\sum_{i}\lambda_{i}}\prod_{i=1}^{n}\exp\left\{-\frac{1}{2}\frac{(\bm{x}_{i}^{\mathrm{\scriptscriptstyle T}}\bm{\beta})^{2}+\lambda_{i}^{2}}{2}\right\}=2^{-\sum_{i}\lambda_{i}}e^{-\sum_{i=1}^{n}\lambda_{i}^{2}/4}\exp\left\{-\frac{1}{2}\left(\frac{\bm{\beta}^{\mathrm{\scriptscriptstyle T}}X^{\mathrm{\scriptscriptstyle T}}X\bm{\beta}}{2}\right)\right\}$
where the last inequality follows from $\cosh(|x|)^{-l}\geq e^{-l|x|}\geq\exp(-x^{2}-l^{2}/4)$ for any $x\in\mathbb{R}$ and $l>0$. Combining together with the remaining parts, denoting $\Sigma^{\star}=(\frac{1}{2}X^{\mathrm{\scriptscriptstyle T}}X+\Sigma_{\beta}^{-1})^{-1}$,
$\int_{(0,\infty)^{n}}p(\bm{\beta}\mid\bm{\kappa},\bm{\lambda})p(\bm{\kappa}\mid\bm{\beta}^{*},\bm{\lambda})$ $\geq(2\pi)^{-p/2}|\Sigma_{\beta}|^{-1/2}\exp\left(-\frac{1}{2}\bm{\beta}^{\mathrm{\scriptscriptstyle T}}\Sigma_{\beta}^{-1}\bm{\beta}-\frac{1}{2}\mathbf{s}^{\mathrm{\scriptscriptstyle T}}\mathbf{s}+\mathbf{s}^{\mathrm{\scriptscriptstyle T}}\Sigma_{\beta}^{-1/2}\bm{\beta}\right)$
×$\quad\times 2^{-\sum_{i}\lambda_{i}}e^{-\sum_{i=1}^{n}\lambda_{i}^{2}/4}\exp\left\{-\frac{1}{2}\left(\frac{\bm{\beta}^{\mathrm{\scriptscriptstyle T}}X^{\mathrm{\scriptscriptstyle T}}X\bm{\beta}}{2}\right)\right\}$
$=2^{-\sum_{i}\lambda_{i}}e^{-\sum_{i=1}^{n}\lambda_{i}^{2}/4}|\Sigma^{\star}|^{1/2}|\Sigma_{\beta}|^{-1/2}$
×$\quad\times\exp\left(-\frac{1}{2}\mathbf{s}^{\mathrm{\scriptscriptstyle T}}(I_{p}-(\Sigma_{\beta}^{1/2}X^{\mathrm{\scriptscriptstyle T}}X\Sigma_{\beta}^{1/2}/2+I_{p})^{-1})\mathbf{s}\right)$
×$\quad\times(2\pi)^{-p/2}|\Sigma^{\star}|^{-1/2}\exp\left(-\frac{1}{2}(\bm{\beta}-\bm{\mu}^{\star})^{\mathrm{\scriptscriptstyle T}}(\Sigma^{\star})^{-1}(\bm{\beta}-\bm{\mu}^{\star})\right)$
$=\delta(\bm{\lambda})N_{p}(\bm{\beta};\bm{\mu}^{\star}(\bm{\lambda}),\Sigma^{\star})$
with $\delta(\bm{\lambda})=2^{-\sum_{i}\lambda_{i}}e^{-\sum_{i=1}^{n}\lambda_{i}^{2}/4}|\Sigma^{\star}|^{1/2}|\Sigma_{\beta}|^{-1/2}\exp\left(-\frac{1}{2}\mathbf{s}^{\mathrm{\scriptscriptstyle T}}(I_{p}-(\Sigma_{\beta}^{1/2}X^{\mathrm{\scriptscriptstyle T}}X\Sigma_{\beta}^{1/2}/2+I_{p})^{-1})\mathbf{s}\right)$ and $\bm{\mu}^{\star}=\bm{\mu}^{\star}(\bm{\lambda})=(\frac{1}{2}X^{\mathrm{\scriptscriptstyle T}}X+\Sigma_{\beta}^{-1})^{-1}\Sigma_{\beta}^{-1/2}\mathbf{s}$. ∎
Finally, to establish uniform ergodicity, we define $q(\bm{\Theta}_{2})\coloneqq q(\bm{\beta})q(\psi)$ in a product form, where each component is
$q(\bm{\beta})=\frac{1}{Z_{\beta}}\min_{\bm{\lambda}\in\{1,\dots,L\}^{n}}N_{p}(\bm{\beta};\bm{\mu}^{\star}(\bm{\lambda}),\Sigma^{\star}),\quad\bm{\beta}\in\mathbb{R}^{p}$
and
$q(\psi)=\frac{1}{Z_{\psi}}\min_{\bm{\lambda}\in\{1,\dots,L\}^{n}}\mathrm{Beta}\Big{(}\psi;a_{\psi}+2n,b_{\psi}-n+\sum_{i=1}^{n}\lambda_{i}\Big{)},\quad\psi\in(0,1)$
where $Z_{\beta}=\int_{\mathbb{R}^{p}}\min_{\bm{\lambda}\in\{1,\dots,L\}^{n}}N_{p}(\bm{\beta};\bm{\mu}^{\star}(\bm{\lambda}),\Sigma^{\star})\mathrm{d}\bm{\beta}<\infty$ and $Z_{\psi}=\int_{0}^{1}\min_{\bm{\lambda}\in\{1,\dots,L\}^{n}}\mathrm{Beta}(\psi;a_{\psi}+2n,b_{\psi}-n+\sum_{i=1}^{n}\lambda_{i})\mathrm{d}\psi<\infty$ are normalizing constants. Then, for any $\bm{\lambda}$, since $N_{p}(\bm{\beta};\bm{\mu}^{\star}(\bm{\lambda}),\Sigma^{\star})\geq Z_{\beta}q(\bm{\beta})$ for all $\bm{\beta}\in\mathbb{R}^{p}$ and $p(\psi\mid\bm{\lambda})\geq Z_{\psi}q(\psi)$ for all $\psi\in(0,1)$ by definition, we have
$k(\bm{\Theta}_{2}\mid\bm{\Theta}_{2}^{*})\geq E_{\bm{\lambda}\sim p(\bm{\lambda}\mid\bm{\beta}^{*})}\{\delta(\bm{\lambda})\}\times Z_{\beta}q(\bm{\beta})Z_{\psi}q(\psi)\geq\min_{\bm{\lambda}}\{\delta(\bm{\lambda})\}Z_{\beta}q(\bm{\beta})Z_{\psi}q(\psi)=\delta^{\star}q(\bm{\Theta}_{2})$
which completes the proof of uniform ergodicity of Algorithm 2, since $\delta^{\star}=\min_{\bm{\lambda}}\{\delta(\bm{\lambda})\}Z_{\beta}Z_{\psi}>0$.
To see $\delta^{\star}>0$ is a constant, defining $\mathbf{s}^{\star}=\Sigma_{\beta}^{1/2}X^{\mathrm{\scriptscriptstyle T}}\tilde{\bm{y}}^{\star}$ where $\tilde{\bm{y}}^{\star}\in\mathbb{R}^{n}$ is a vector with $i$th element $L(y_{i}-1/2)$, which does not depend on $\bm{\lambda}$, we have
$\min_{\bm{\lambda}}\{\delta(\bm{\lambda})\}=2^{-nL}e^{-nL^{2}/4}|\Sigma^{\star}|^{1/2}|\Sigma_{\beta}|^{-1/2}\exp\left(-\frac{1}{2}(\mathbf{s}^{\star})^{\mathrm{\scriptscriptstyle T}}(I_{p}-(\Sigma_{\beta}^{1/2}X^{\mathrm{\scriptscriptstyle T}}X\Sigma_{\beta}^{1/2}/2+I_{p})^{-1})\mathbf{s}^{\star}\right)$
since $(I_{p}-(\Sigma_{\beta}^{1/2}X^{\mathrm{\scriptscriptstyle T}}X\Sigma_{\beta}^{1/2}/2+I_{p})^{-1})$ is nonnegative definite by Lemma S.1.
The uniform ergodicity of Algorithm 1 for cobin regression is based on the simpler Markov transition density with $\bm{\Theta}_{2}=\bm{\beta}$,
$k(\bm{\beta}\mid\bm{\beta}^{*})$ $=\sum_{\lambda=1}^{L}\left\{\int_{(0,\infty)^{n}}p(\bm{\beta}\mid\bm{\kappa},\lambda)p(\bm{\kappa}\mid\bm{\beta}^{*},\lambda)\mathrm{d}\bm{\kappa}\right\}p(\lambda\mid\bm{\beta}^{*})$
$\geq\sum_{\lambda=1}^{L}\delta(\lambda)N_{p}(\bm{\beta};\bm{\mu}^{\star}(\lambda),\Sigma^{\star})p(\lambda\mid\bm{\beta}^{*})$
$\geq\min_{\lambda}\{\delta(\lambda)\}Z_{\beta}q(\bm{\beta})=\delta^{\star}q(\bm{\beta})$
where previous vector $\bm{\lambda}$ inputs are now corresponding to $\lambda\bm{1}_{n}$, and previous minimum over $\{1,\dots,L\}^{n}$ is now corresponding to minimum over $\{1,\dots,L\}$. This completes the uniform ergodicity of the cobin regression blocked Gibbs sampler.
S.1.3 Proof of Proposition 2
We recall from Feller, (1948) and Devroye, (1986) §5.6 that the Kolmogorov distribution $\mathcal{K}$ admits two different density representations:
$p_{\textsc{k}}(x)=8\sum_{n=0}^{\infty}(-1)^{n}(n+1)^{2}x\exp\{-2(n+1)^{2}x^{2}\}$ (S.3)
and
$p_{\textsc{k}}(x)=\frac{(2\pi)^{1/2}}{x}\sum_{n=0}^{\infty}\left\{\frac{(2n+1)^{2}\pi^{2}}{4x^{3}}-\frac{1}{x}\right\}\exp\left\{-\frac{(2n+1)^{2}\pi^{2}}{8x^{2}}\right\}$ (S.4)
It is well known that the squared Kolmogorov random variable can be represented as an infinite convolution of exponential random variables, i.e. $\mathcal{K}^{2}\stackrel{{\scriptstyle d}}{{=}}0.5\sum_{k=1}^{\infty}\epsilon_{k}/k^{2}$, $\epsilon_{k}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}\mathrm{Exp}(1)$ (Andrews and Mallows,, 1974, §4). By definition of Kolmogorov-Gamma, we have $\mathrm{KG}(1,0)\stackrel{{\scriptstyle d}}{{=}}\mathcal{K}^{2}/\pi^{2}$. Thus, applying change of variables with both density representations,
${p_{\textsc{kg}}}(x;1,0)=\sum_{n=0}^{\infty}(-1)^{n}4\pi^{2}(n+1)^{2}\exp\left\{-2\pi^{2}(n+1)^{2}x\right\}$ (S.5)
and
${p_{\textsc{kg}}}(x;1,0)=\frac{1}{(2\pi)^{1/2}}\sum_{n=0}^{\infty}\left\{\frac{(2n+1)^{2}}{4x^{5/2}}-\frac{1}{x^{3/2}}\right\}\exp\left\{-\frac{(2n+1)^{2}}{8x}\right\}.$ (S.6)
It is easy to see how the first sum terms simplify into the $a_{n}^{R}(x)$ terms given in the statement of Proposition 2. Obtaining the second representation requires re-indexing the sum so that the even parts correspond to the term with $4x^{5/2}$ in the denominator and the odd parts correspond to the term with $x^{3/2}$ in the denominator. After that, the form for the $a^{L}_{n}(x)$ terms given in Proposition 2 emerges.
S.1.4 Proof of Lemma 1
It suffices to show that both $a_{n}^{L}(x)$ and $a_{n}^{R}(x)$ are decreasing in $n$ in the indicated interval. We first consider $a_{n}^{R}(x)$, which we note to be decreasing in $n$ if $(n+1)^{2}\exp(-2\pi^{2}(n+1)^{2}x)<n^{2}\exp(-2\pi^{2}n^{2}x)$. But this is equivalent to inequality $(\log(n+1)-\log(n))/((2n+1)\pi^{2})<x$. Since the numerator of the left term decreases in $n$ and the denominator increases in $n$, the inequality is satisfied for all $n$ if it holds for $n=1$. But that reduces to the inequality $\log(2)/(3\pi^{2})<x$, and we conclude that $a_{n}^{R}(x)$ is decreasing in $n$ within $(\log(2)/3\pi^{2},\infty)$.
We now show that $a_{n}^{L}(x)$ decreases in $n$ for $x<1/4$. If $n$ is even, $a_{n}^{L}(x)-a_{n+1}^{L}(x)$ can be expressed as a positive number times $(n+1)^{2}-4x$, which is positive if $x<1/4$. On the other hand, if $n$ is odd, $\log a_{n}^{L}(x)-\log a_{n+1}^{L}(x)>0$ if $(n+1)/(2x)+\log x-2\log(n+2)+2\log 2>0$. We note that the derivative of this expression with respect to $x$ is $1/x-(n+1)/(2x^{2})$, which is negative for $x\in(0,(n+1)/2)$, and so the difference is decreasing on the interval $(0,(n+1)/2)$ that includes $(0,1/4)$. Finally, we observe that the inequality $(n+1)/(2x)+\log x-2\log(n+2)+2\log 2>0$ holds for $x=1/4$, and therefore $a_{n}^{L}(x)$ is decreasing in $n$ for all $x\in(0,1/4)$.
Therefore, so long as $t\in(\log(2)/(3\pi^{2}),1/4)$, the sequence $a_{n}(x;c,t)$ is monotonically decreasing in $n$ regardless of $c$.
S.1.5 Proof of Proposition 3
The proof of this proposition has three parts. First, we show that the optimal cutoff $t^{*}$ is independent of $c$. Next, we provide an upper bound of the expected number of outer loop iterations for any $c$, i.e. the uniform lower bound on the acceptance probability of a proposal. Finally, we provide the uniform upper bound of the expected number of inner loop iterations.
[…]
$a_{0}(x;c,t)=\begin{cases}\frac{\sinh(c/2)}{c/2}\exp(-c^{2}x/2)\frac{\exp(-\frac{1}{8x})}{\pi^{1/2}(2x)^{5/2}},&0<x<t,\\ \frac{\sinh(c/2)}{c/2}\exp(-c^{2}x/2)4\pi^{2}\exp(-2\pi^{2}x)&t\leq x\end{cases}$
We denote the c.d.f. of $\mathrm{GIG}(-3/2,c^{2},1/4)$ as $\texttt{pgig}(t\mid p=-3/2,a=c^{2},b=1/4)$ and the c.d.f. of inverse gamma distribution with parameter $3/2$ and $1/8$ (with density proportional to $x^{-5/2}e^{-1/(8x)}$) as $\texttt{pgig}(t\mid p=-3/2,a=0,b=1/4)$ to unify the notation. Then,
$A^{L}(c,t)$ $=\int_{0}^{t}a_{0}(x;c,t)\mathrm{d}x$
$=\int_{0}^{t}\frac{\sinh(c/2)}{c/2}\exp(-c^{2}x/2)\frac{\exp(-\frac{1}{8x})}{\pi^{1/2}(2x)^{5/2}}\mathrm{d}x$ (S.7)
$=\frac{\sinh(c/2)}{c/2}\frac{Z(c)}{\pi^{1/2}2^{5/2}}\int_{0}^{t}\frac{1}{Z(c)}x^{-5/2}\exp\left(-\frac{1}{2}\left(c^{2}x+\frac{1}{4x}\right)\right)\mathrm{d}x,$
$=\frac{\sinh(c/2)}{c/2}(|c|+2)\exp(-|c|/2)\times\texttt{pgig}(t\mid p=-3/2,a=c^{2},b=1/4)$ (S.8)
where $Z(c)=\frac{2K_{-3/2}(|c|/2)}{(2|c|)^{-3/2}}$ if $c\neq 0$ or $Z(c)=\pi^{1/2}2^{7/2}$ if $c=0$ is a normalizing constant, and $K_{-3/2}$ is a modified Bessel function of the second kind. For the case when $c\neq 0$, we used $K_{-3/2}(|c|/2)=(\pi/|c|)^{1/2}\exp(-|c|/2)(1+2/|c|)$ in simplification. Also,
$A^{R}(c,t)$ $=\int_{t}^{\infty}a_{0}(x;c,t)\mathrm{d}x=\frac{\sinh(c/2)}{c/2}\int_{t}^{\infty}\exp(-c^{2}x/2)4\pi^{2}\exp(-2\pi^{2}x)\mathrm{d}x$ (S.9)
$=\frac{\sinh(c/2)}{c/2}\frac{4\pi^{2}}{2\pi^{2}+c^{2}/2}\exp\left\{-(2\pi^{2}+c^{2}/2)t\right\}.$ (S.10)
We now show that the optimal cutoff $t^{*}$ that minimizes $A^{L}(c,t)+A^{R}(c,t)$ does not depend on $c$. For a given $c$, $A^{L}(c,t)$ and $A^{R}(c,t)$ are both differentiable in $t$. From expressions (S.7) and (S.9), the optimal cutoff $t^{*}$ minimizes
$\int_{0}^{t}\frac{1}{\pi^{1/2}(2x)^{5/2}}\exp\left(-\frac{1}{2}\left(c^{2}x+\frac{1}{4x}\right)\right)\mathrm{d}x+\int_{t}^{\infty}4\pi^{2}\exp\left\{-\left(\frac{c^{2}}{2}+2\pi^{2}\right)x\right\}\mathrm{d}x$ (S.11)
as a function of $t$, since the common $\sinh(c/2)/(c/2)$ term is never zero and does not depend on $t$. We claim that there is a unique $t^{*}\approx 0.050239$ within the interval $(\log(2)/(3\pi^{2}),1/4)$ that minimizes (S.11). To see this, differentiating (S.11) with respect to $t$ and equating to zero, $t^{*}$ is a solution of
$\frac{1}{\pi^{1/2}(2t)^{5/2}}\exp\left(-\frac{1}{8t}\right)-4\pi^{2}\exp\left(-2\pi^{2}t\right)=0$ (S.12)
after canceling the $\exp(-c^{2}t/2)$ term, so the minimizer does not depend on $c$. Rearranging terms,
$12^{9/2}\pi^{5/2}\exp\left(-2\pi^{2}t+\frac{1}{8t}+\frac{5}{2}\log t\right)=$ 1 (S.13)
whose solution is $t^{*}\approx 0.050239$ which lies within the bounds specified by Lemma 1. Since the derivative of the LHS of (S.13) is always negative on $t>0$, $t^{*}$ is unique.
Part 2. It can be easily checked that for fixed $t$, $A^{L}(c,t)$ and $A^{R}(c,t)$ are both continuous in $c$. Also, we have $A^{L}(0,t^{*})+A^{R}(0,t^{*})\approx 1.089002$. We claim that for fixed $t$, $A^{L}(c,t)\to 1$ and $A^{R}(c,t)\to 0$ as $c\to\infty$ to ensure that $A^{L}(c,t^{*})+A^{R}(c,t^{*})$ converges to 1 as $c\to\infty$ and thus the sampler is not ill behaved in the large $c$ regime. To see $\lim_{c\to\infty}A^{L}(c,t)=1$, denoting $G\sim\mathrm{GIG}(-3/2,c^{2},1/4)$, we have $1-\texttt{pgig}(t\mid p=-3/2,a=c^{2},b=1/4)\leq E(G)/t=1/\{2t(c+2)\}$ by the Markov inequality, thus
$1-1/(2tc+4t)\leq\texttt{pgig}(t\mid p=-3/2,a=c^{2},b=1/4)\leq 1$
Combining with $\lim_{c\to\infty}\frac{\sinh(c/2)}{c/2}(|c|+2)e^{-|c|/2}=1$ for the LHS of the inequality, we have $\lim_{c\to\infty}A^{L}(c,t)=1$. The fact that $\lim_{c\to\infty}A^{R}(c,t)=0$ is easily deduced from (S.10).
For the optimal choice of $t^{*}\approx 0.050239$, numerical investigation shows that $A^{L}(c,t^{*})+A^{R}(c,t^{*})$ attains a maximum value of approximately 1.145583 as a function of $c$ when $c=10.134$. Therefore, $M$ is bounded above by 1.1456 for any given $c$. Hence, the average probability of accepting a proposal is uniformly bounded below by 0.8729.
Part 3. We follow Proposition 3 of Polson et al., (2013), where the probability of deciding to accept or reject a proposal $X$ upon checking the $m$th partial sum $S_{m}(X)$ is given by
$\frac{1}{A^{L}(c,t^{*})+A^{R}(c,t^{*})}\int_{0}^{\infty}\left\{a_{m-1}(x;c,t^{*})-a_{m}(x;c,t^{*})\right\}\mathrm{d}x$
The first few of these probabilities for the worst possible envelope (that is, when $c=10.34$) are presented below.
Probability of deciding to accept or reject upon computing $m^{\text{th}}$ series term
$m$ $1$ $2$ $3$ $4$ $5$ $6$
Prob. $0.127$ $2.068\times 10^{-4}$ $2.226\times 10^{-7}$ $3.124\times 10^{-11}$ $5.894\times 10^{-16}$ $1.513\times 10^{-21}$
We see that the probabilities are decaying quickly, guaranteeing that with very high probability, only a small handful of $S_{m}(x)$ terms will need to be computed. From the above, using the worst-case $c=10.134$, the expected number of series terms that will need to be computed to decide whether to accept or reject a proposal is 1.1274624.
S.2 Pseudocode for KG(1,c) sampler
Algorithm S.1 Sampling from $\text{KG}(1,c)$
1:Input: Parameter $c$, cutoff value $t\in(0.0234,0.25)$ (optimal $t^{*}=0.050239$)
2:$\tilde{A}^{L}\leftarrow(|c|+2)\exp(-|c|/2)\texttt{pGIG}(t\mid p=-3/2,a=c^{2},b=1/4)$triangleright proportional to $A^{L}(c,t)$
3:$\tilde{A}^{R}\leftarrow 4\pi^{2}\exp\{-(2\pi^{2}+c^{2}/2)t\}/(2\pi^{2}+c^{2}/2)$triangleright proportional to $A^{R}(c,t)$
4:repeat
5:     Generate $U,V\sim U(0,1)$
6:     if $U<\tilde{A}^{R}/(\tilde{A}^{L}+\tilde{A}^{R})$ then $\triangleright$ $\sinh(c/2)/(c/2)$ are canceled out thus not calculated
7:         $X\leftarrow t+E/(2\pi^{2}+c^{2}/2)$ where $E\sim\text{Exp}(1)$triangleright Truncated exponential
8:     else
9:         $\triangleright$ Truncated GIG
10:              $X\sim\mathrm{GIG}(p=-3/2,a=c^{2},b=1/4)$triangleright If $c=0$, $X\sim\mathrm{InvGamma}(3/2,1/8)$
11:         until $X<t$
12:     end if
13:     $S\leftarrow a_{0}(X)$, $Y\leftarrow VS$, $m\leftarrow 0$
14:     repeat
15:         $m\leftarrow m+1$
16:         if $m$ is odd then
17:              $S\leftarrow S-a_{m}(X)$; if $Y<S$, then return $m$
18:         else
19:              $S\leftarrow S+a_{m}(X)$; if $Y>S$, then break
20:         end if
21:     until FALSE
22:until FALSE
The Algorithm S.1 describes the pseudocode for the KG$(1,c)$ sampler using the alternating series method. For the part involving the sampling and evaluation of the c.d.f of the GIG distribution with half-integer parameter $p=-3/2$, we employ the result of Peña and Jauch, (2025). When $c=0$, sampling from $\mathrm{GIG}(-3/2,c^{2},1/4)$ in line 10 is replaced with $\mathrm{InvGamma}(3/2,1/8)$, and c.d.f. evaluation in line 2 is replaced with c.d.f. of $\mathrm{InvGamma}(3/2,1/8)$. For the choice of optimal $t^{*}=0.050239$, we have $\texttt{pGIG}(t^{*}\mid p=-3/2,a=0,b=1/4)=0.1735472$.
To sample from the truncated GIG distribution supported on $(0,t)$, we use a simple rejection method by drawing a GIG variate until it falls in $(0,t)$. Letting $t$ be fixed at the optimal value $t^{*}=0.050239$, the expected number of draws depending on the choice of $c$ is $1/\texttt{pGIG}(t^{*}\mid p=-3/2,a=c^{2},b=1/4)$. By comparing the integrand, it can be verified that $\texttt{pGIG}(t^{*}\mid p=-3/2,a,b=1/4)$ is an increasing function of $a$ and achieves a minimum value of $0.1735472$ at $a=0$. Thus, the expected number of GIG draws to obtain a single truncated GIG sample is not larger than $1/0.1735472\approx 5.7622$. In practice, we found that this simple rejection method is much more efficient and numerically stable than transforming a uniform random variable through the inverse of the c.d.f. of a truncated GIG distribution.
S.3 Details of cobin and micobin distributions
S.3.1 Cobin distribution as an exponential dispersion model
Following Jørgensen, (1986, §2), we describe in detail how the proposed continuous binomial distribution arises from an exponential dispersion model generated from the uniform distribution. Let $Q$ be a probability measure that corresponds to a uniform distribution. Then $M(s)=\int\exp(sy)\mathrm{d}Q(x)=(e^{s}-1)/s=\exp(B(s))$ for $s\in\mathbb{R}$.
First, consider the set of $\lambda$ values such that $M(s)^{\lambda}=\{(e^{s}-1)/s)\}^{\lambda}$ is a moment generating function of some distribution $Q_{\lambda}$. This set corresponds to the non-negative integers $\{\lambda\in\mathbb{R}\backslash\{0\}:M^{\lambda}$ is the m.g.f. of some distribution $Q_{\lambda}\}=\{1,2,\dots,\}$. To see this, from Theorem 3.1 (vi) of Jørgensen, (1986), it suffices to show that the analytic continuation of $M$ to the complex plane has at least one simple zero, which is confirmed by checking $M(z)=0$ has solutions $z=\pm 2\pi i,\pm 4\pi i,\dots$, where $i$ is an imaginary unit.
Then, the corresponding probability measure $Q_{\lambda}$ is the $\lambda$-fold i.i.d. convolution of uniform distribution, i.e., the Irwin-Hall distribution defined on the interval $(0,\lambda)$. Based on $Q_{\lambda}$ ($\lambda=1,2,\dots$), we consider the probability measure $Q_{\lambda,\theta}$ that satisfies
$\frac{\mathrm{d}Q_{\lambda,\theta}}{\mathrm{d}Q_{\lambda}}=\exp\{x\theta-\lambda B(\theta)\},$
where the left-hand side denotes the Radon–Nikodym derivative. Hence, $Q_{\lambda,\theta}$ is the distribution obtained from exponentially tilting $Q_{\lambda}$ by $e^{x\theta}$. Transforming to $y=x/\lambda$, we obtain the exponential dispersion model $P_{\lambda,\theta}$ that satisfies $\frac{\mathrm{d}P_{\lambda,\theta}}{\mathrm{d}P_{\lambda}}=\exp\{\lambda y\theta-\lambda B(\theta)\}.$ Since the density corresponding to $P_{\lambda}$ (Irwin-Hall scaled by $1/\lambda$) is $h(y,\lambda)=\frac{\lambda}{(\lambda-1)!}\sum_{k=0}^{\lambda}(-1)^{k}{\lambda\choose k}\max\{(\lambda y-k),0\}^{\lambda-1},$ the density corresponding to $P_{\lambda,\theta}$ coincides with (1).
Table S.1 presents a comparative summary with normal, gamma, and inverse Gaussian exponential dispersion families.
Table S.1: Illustration of the derivation of cobin distribution as a continuous exponential dispersion model, along with a comparison with normal, gamma, and inverse Gaussian distributions.
    exponential tilting by $\theta$   $\lambda$-fold convolution and scale by $\lambda^{-1}$
$\xrightarrow{\qquad\quad}$         $\xrightarrow{\qquad\quad}$
$H\sim\mathrm{Unif}(0,1)$ $Y\sim\mathrm{cobin}(\theta,1)$ $Y\sim\mathrm{cobin}(\theta,\lambda^{-1})$
$h(y)=1$ $\exp(\theta y-B(\theta))h(y)$ $\exp(\lambda\theta y-\lambda B(\theta))h(y,\lambda)$
$0\leq y\leq 1$ $\theta\in\mathbb{R}$ $\lambda\in\mathbb{N}$
$H_{1}\sim N(0,1)$ $Y\sim N(\theta,1)$ $Y\sim N(\theta,\lambda^{-1})$
$h_{1}(y)=(2\pi)^{-1/2}\exp(-y^{2}/2)$ $\exp(\theta y-B_{1}(\theta))h_{1}(y)$ $\exp(\lambda\theta y-\lambda B_{1}(\theta))h_{1}(y,\lambda)$
$y\in\mathbb{R}$ $\theta\in\mathbb{R}$ $\lambda>0$
$H_{2}\sim\mathrm{Exp}(1)$ $Y\sim\mathrm{Exp}(1-\theta)$ $Y\sim\mathrm{Gamma}(\lambda,\lambda(1-\theta))$
$h_{2}(y)=\exp(-y)$ $\exp(\theta y-B_{2}(\theta))h_{2}(y)$ $\exp(\lambda\theta y-\lambda B_{2}(\theta))h_{2}(y,\lambda)$
$y\geq 0$ $\theta<1$ $\lambda>0$
$H_{3}\sim\mathrm{InvGamma}(1/2,1/2)$ $Y\sim\mathrm{InvGau}((-2\theta)^{-1/2},1)$ $Y\sim\mathrm{InvGau}((-2\theta)^{-1/2},\lambda)$
$h_{3}(y)=(2\pi y^{3})^{-1/2}\exp(-\frac{1}{2y})$ $\exp(\theta y-B_{3}(\theta))h_{3}(y)$ $\exp(\lambda\theta y-\lambda B_{3}(\theta))h_{3}(y,\lambda)$
$y>0$ $\theta<0$ $\lambda>0$
$B(\theta)=\log E(e^{\theta H})=\log\left(\frac{e^{\theta}-1}{\theta}\right),\quad h(y,\lambda)=\text{density of }\frac{1}{\lambda}\sum_{l=1}^{\lambda}H^{(l)}$
$B_{1}(\theta)=\log E(e^{\theta H_{1}})=\frac{\theta^{2}}{2},\quad h_{1}(y,\lambda)=\text{density of }\frac{1}{\lambda}\sum_{l=1}^{\lambda}H_{1}^{(l)}=\frac{\exp(-\lambda y^{2}/2)}{\sqrt{2\pi/\lambda}}$
$B_{2}(\theta)=\log E(e^{\theta H_{2}})=-\log(1-\theta),\quad h_{2}(y,\lambda)=\text{density of }\frac{1}{\lambda}\sum_{l=1}^{\lambda}H_{2}^{(l)}=\frac{\lambda^{\lambda}y^{\lambda-1}\exp(-\lambda y)}{\Gamma(\lambda)}$
$B_{3}(\theta)=\log E(e^{\theta H_{3}})=-(-2\theta)^{1/2},\quad h_{3}(y,\lambda)=\text{density of }\frac{1}{\lambda}\sum_{l=1}^{\lambda}H_{3}^{(l)}=\frac{\lambda^{1/2}\exp(-\lambda/(2y))}{\sqrt{2\pi y^{3}}}$
S.3.2 Further details of cobin and micobin distributions
The density of the micobin distribution is
$p_{\mathrm{micobin}}(y;\theta,\psi)=\sum_{\lambda=1}^{\infty}\lambda(1-\psi)^{\lambda-1}\psi^{2}h(y,\lambda)\frac{e^{\lambda\theta y}}{\{(e^{\theta}-1)/\theta\}^{\lambda}\}},\quad 0\leq y\leq 1$ (S.14)
which follows from its definition as an hierarchical model $Y\mid\lambda\sim\mathrm{cobin}(\theta,\lambda^{-1})$, $(\lambda-1)\sim\mathrm{negbin}(2,\psi)$. $p_{\mathrm{micobin}}$ is continuous in $y\in[0,1]$, since mixture component densities $p_{\mathrm{cobin}}(y;\theta,\lambda^{-1})$ are all continuous. When $\psi\to 1$, it reduces to $\mathrm{cobin}(\theta,1)$. It has nonzero density at boundary values,
$p_{\mathrm{micobin}}(0;\theta,\psi)=\psi^{2}p_{\mathrm{cobin}}(0;\theta,1)=\psi^{2}\frac{\theta}{e^{\theta}-1}$
$p_{\mathrm{micobin}}(1;\theta,\psi)=\psi^{2}p_{\mathrm{cobin}}(1;\theta,1)=\psi^{2}\frac{\theta e^{\theta}}{e^{\theta}-1}$
Unlike cobin, which belongs to the exponential dispersion family and its likelihood is guaranteed to be log-concave in terms of $\theta$, there is no guarantee that the likelihood of micobin distribution is log-concave.
The c.d.f. of $\mathrm{cobin}(\theta,\lambda^{-1})$ can be obtained by integrating $p_{\mathrm{cobin}}$ term-by-term. Its form when $\lambda=1$ is available from Loaiza-Ganem and Cunningham, (2019), which is $F_{\mathrm{cobin}}(z;\theta,1)=(e^{\theta z}-1)/(e^{\theta}-1)$. Assuming $\lambda\geq 2$,
$F_{\mathrm{cobin}}(z;\theta,\lambda^{-1})$ $=\frac{\lambda}{(\lambda-1)!}\sum_{k=0}^{\lambda}(-1)^{k}{\lambda\choose k}\exp(-\lambda B(\theta))\int_{0}^{z}\max\{(\lambda y-k),0\}^{\lambda-1}\exp(\lambda y\theta)\mathrm{d}y$
$=\frac{\lambda}{(\lambda-1)!}\sum_{k=0}^{\lambda}(-1)^{k}{\lambda\choose k}\exp(-\lambda B(\theta))\int_{k/\lambda}^{z}(\lambda y-k)^{\lambda-1}\exp(\lambda y\theta)\mathrm{d}y$
where we used $\max\{(\lambda y-k),0\}=0$ if $y\leq k/\lambda$. The integral term becomes
$\int_{k/\lambda}^{z}(\lambda y-k)^{\lambda-1}\exp(\lambda y\theta)\mathrm{d}y=\begin{cases}\lambda^{-2}(\lambda z-k)^{\lambda}&\theta=0\\ \frac{e^{\theta k}}{\lambda(-\theta)^{\lambda}}\gamma(\lambda,-\theta(\lambda z-k))&\theta\neq 0\end{cases}$
where $\gamma(\lambda,x)=\int_{0}^{x}t^{\lambda-1}e^{-t}\mathrm{d}x$ is a lower incomplete gamma function. In practice, many existing software programs only support the calculation of the lower incomplete gamma function with positive $x$, i.e., when $\theta$ is negative. To resolve this, one can use symmetry of cobin distributions between $\mathrm{cobin}(\theta,\lambda)$ and $\mathrm{cobin}(-\theta,\lambda)$, which yields $F_{\mathrm{cobin}}(z;\theta,\lambda^{-1})=1-F_{\mathrm{cobin}}(1-z;-\theta,\lambda^{-1})$.
The c.d.f. of $\mathrm{micobin}(\theta,\psi)$ is simply a weighted sum of cobin c.d.f.s,
$F_{\mathrm{micobin}}(z;\theta,\psi)=\sum_{\lambda=1}^{\infty}\lambda(1-\psi)^{\lambda-1}\psi^{2}F_{\mathrm{cobin}}(z;\theta,\lambda^{-1}).$
The random variate generation of $\mathrm{cobin}(\theta,\lambda^{-1})$ can be easily done by taking an average of $\lambda$ i.i.d. $\mathrm{cobin}(\theta,1)$ variables. The sampling from $\mathrm{cobin}(\theta,1)$ can be done by $F_{\mathrm{cobin}}^{-1}(U;\theta,1)$ with $U\sim\mathrm{Unif}(0,1)$, where $F_{\mathrm{cobin}}^{-1}(u;\theta,1)=\theta^{-1}\log(ue^{\theta}-u+1)$ (Loaiza-Ganem and Cunningham,, 2019). Note that if $\theta=0$ then $\mathrm{cobin}(0,1)$ is a uniform distribution. Random variate generation of micobin directly follows from its definition.
S.3.3 Cobit link function and variance function
The cobit link function $g_{\mathrm{cobit}}:(0,1)\to\mathbb{R}$ has an inverse $g^{-1}_{\mathrm{cobit}}(x)=B^{\prime}(x)=e^{\theta}/(e^{\theta}-1)-\theta^{-1}$. Although the expression for $g_{\mathrm{cobit}}$ is not available analytically, the numerical inversion of $B^{\prime}(x)$ can be easily done with the Newton–Raphson algorithm. Under cobin regression with canonical link, the sufficient statistic of $\bm{\beta}$ is $X^{\mathrm{\scriptscriptstyle T}}\bm{y}$, and the log-likelihood is guaranteed to be concave (Agresti,, 2015). The cobit link function satisfies $\lim_{x\to-\infty}g^{-1}_{\mathrm{cobit}}(\pi x)/g_{\mathrm{cauchit}}^{-1}(x)=\lim_{x\to+\infty}\{1-g^{-1}_{\mathrm{cobit}}(\pi x)\}/\{1-g_{\mathrm{cauchit}}^{-1}(x)\}=1$, where $g_{\mathrm{cauchit}}^{-1}=\arctan(x)/\pi+0.5$ (Koenker and Yoon,, 2009) is an inverse of the cauchit link function. Thus, up to scale difference in the linear predictor, the cobit link maps a large linear predictor to the mean around 0 or 1, asymptotically at the same rate as the cauchit link. Compared to the logit link, this significantly reduces the influence of large outlying predictors, also contributing to robustness; see Gelman and Hill, (2007, §6.6) for a discussion. See Figure S.5 for a comparison between cauchit and logit link functions.
Figure S.5: Comparison of link and variance functions. (Left) Inverse of cobit link $g^{-1}_{\mathrm{cobit}}(x)=B^{\prime}(x)$, Cauchit link and logit link; cauchit and logit are scaled such that derivatives at zero are the same. (Right) Variance function $V(\mu)=B^{\prime\prime}\{(B^{\prime})^{-1}(\mu)\}$ with variance function of beta regression model scaled by $1/3$.
We argue that there is no reason that logit link is preferred over other link functions when dealing with proportions that are not interpreted as a probability. Even if it is interpreted as a probability, logit link does not give the exactly same interpretation as the logistic regression in terms of the odds ratio, and extra care is required. For example, if $y$ is the area of dense breast tissue divided by the whole breast area on a mammogram (Peplonska et al.,, 2012), $y$ is not a probability that a patient has breast cancer. Even if $y$ is a probability of some event, the corresponding regression model (either cobin or beta) yields
$\log\left\{\frac{E(y\mid\bm{x})}{1-E(y\mid\bm{x})}\right\}=\bm{x}^{\mathrm{\scriptscriptstyle T}}\bm{\beta}\neq E\left\{\log\left(\frac{y}{1-y}\right)\mid\bm{x}\right\}\neq\log E\left\{\left(\frac{y}{1-y}\right)\mid\bm{x}\right\}$
thus the “log odds” should be newly defined as a quantity $\log\frac{E(y\mid\bm{x})}{1-E(y\mid\bm{x})}$, which is different from the expected log odds, and also different from the log of expected odds.
The variance function $V(\mu)=B^{\prime\prime}\{(B^{\prime})^{-1}(\mu)\}$ has a maximum of $1/12$ at $\mu=1/2$. This is in constrast to the “variance function” of beta regression $\mu(1-\mu)$, which has a maximum of $1/4$ at $\mu=1/2$ and satisfies $\operatorname*{var}(Y)=\mu(1-\mu)/(1+\phi)$ for $Y\sim\mathrm{Beta}(\mu,\phi)$ (mean, precision parametrization). The range of the variance function of cobin is less than that of the beta, as the beta density spikes at the boundaries for small $\phi$, leading to a higher variance. Another notable difference is the behavior of the variance function when $\mu$ is close to zero or one, where $\mu(1-\mu)$ approaches to 0 rapidly whereas $V(\mu)=B^{\prime\prime}\{(B^{\prime})^{-1}(\mu)\}$ approaches to 0 smoothly. See the right panel of Figure S.5 for a comparison.
S.4 Detailed derivation of Gibbs samplers and EM algorithm
S.4.1 Fixed effects models
We first add details on the derivations of Algorithm 1 and Algorithm 2, where we suppressed the notations conditioning on data $\bm{y}$ for conciseness. The update for $\lambda$ (or $\lambda_{i}$ for micobin) is based on the posterior under the non-augmented model,
$p(\bm{\beta},\lambda\mid\bm{y})$ $\propto p(\bm{\beta})p_{\lambda}(\lambda)\prod_{i=1}^{n}p_{\mathrm{cobin}}(y_{i};\bm{x}_{i}^{\mathrm{\scriptscriptstyle T}}\bm{\beta},\lambda^{-1})$ (cobin)
$p(\bm{\beta},\bm{\lambda}\mid\psi,\bm{y})$ $\propto p(\bm{\beta})\left\{\prod_{i=1}^{n}\lambda_{i}(1-\psi)^{\lambda_{i}-1}\psi^{2}\right\}\left\{\prod_{i=1}^{n}p_{\mathrm{cobin}}(y_{i};\bm{x}_{i}^{\mathrm{\scriptscriptstyle T}}\bm{\beta},\lambda_{i}^{-1})\right\}$ (micobin)
which yields step 1 of both algorithms by conditioning on $\bm{\beta}$. Given $\lambda$ (or $\bm{\lambda}$ for micobin), the conditional posterior under the augmented model is, combining with (11),
$p(\bm{\beta},\bm{\kappa}\mid\lambda,\bm{y})$ $\propto p(\bm{\beta})\prod_{i=1}^{n}\exp\{{\lambda(y_{i}-0.5)\bm{x}_{i}^{\mathrm{\scriptscriptstyle T}}\bm{\beta}}-\kappa_{i}(\bm{x}_{i}^{\mathrm{\scriptscriptstyle T}}\bm{\beta})^{2}/2\}{p_{\textsc{kg}}}(\kappa_{i};\lambda,0),$ (cobin)
$p(\bm{\beta},\bm{\kappa}\mid\bm{\lambda},\bm{y})$ $\propto p(\bm{\beta})\prod_{i=1}^{n}\exp\{{\lambda_{i}(y_{i}-0.5)\bm{x}_{i}^{\mathrm{\scriptscriptstyle T}}\bm{\beta}}-\kappa_{i}(\bm{x}_{i}^{\mathrm{\scriptscriptstyle T}}\bm{\beta})^{2}/2\}{p_{\textsc{kg}}}(\kappa_{i};\lambda_{i},0).$ (micobin)
From Theorem 2, both algorithms’ step 2 of updating KG random variables follows by conditioning on $\bm{\beta}$. Conditioning on $\bm{\kappa}$, with normal prior $\bm{\beta}\sim N(\bm{0},\Sigma_{\beta})$, step 3 follows by observing that $\exp\{{\lambda(y_{i}-0.5)\bm{x}_{i}^{\mathrm{\scriptscriptstyle T}}\bm{\beta}}-\kappa_{i}(\bm{x}_{i}^{\mathrm{\scriptscriptstyle T}}\bm{\beta})^{2}/2\}$ is proportional to $N(\lambda(y_{i}-0.5)/\kappa_{i};\bm{x}_{i}^{\mathrm{\scriptscriptstyle T}}\bm{\beta},\kappa_{i}^{-1})$ in $\bm{\beta}$, where $N(y;\mu,\sigma^{2})$ denotes the $N(\mu,\sigma^{2})$ density evaluated at $y$. Finally, step 4 of Algorithm 2 comes from the fact that the contribution from the latent negative binomial model to the likelihood of $\psi$ is $\prod_{i=1}^{n}\{\lambda_{i}(1-\psi)^{\lambda_{i}-1}\psi^{2}\}$.
S.4.2 Mixed effect models
We first derive a generic partially collapsed Gibbs sampler (Van Dyk and Park,, 2008) for cobin and micobin mixed effect models, where fixed effect and random effect coefficients are jointly updated. This is made possible due to KG augmentation, leveraging conditional normal likelihood and normal-normal conjugacy. Then we specialize in random intercept and spatial regression models. Consider the following general mixed model setting with random effect $\bm{u}\in\mathbb{R}^{q}$,
$g_{\mathrm{cobit}}(E\{y_{i}\mid u_{i},\bm{x},\bm{z}\})=\bm{x}_{i}^{\mathrm{\scriptscriptstyle T}}\bm{\beta}+\bm{z}_{i}^{\mathrm{\scriptscriptstyle T}}\bm{u},\quad\bm{u}\sim N_{q}(\bm{0},\Sigma_{u}(\vartheta)),\quad i=1,\dots,n,$ (S.15)
along the cobin or micobin response distribution with parameter $\theta_{i}=\eta_{i}=\bm{x}_{i}^{\mathrm{\scriptscriptstyle T}}\bm{\beta}+\bm{z}_{i}^{\mathrm{\scriptscriptstyle T}}\bm{u}$. Here $\vartheta$ is random effect covariance parameter(s), $\bm{z}_{i}\in\mathbb{R}^{q}$ corresponds to the random effect covariate, and $Z=(\bm{z}_{1}^{\mathrm{\scriptscriptstyle T}},\dots,\bm{z}_{n}^{\mathrm{\scriptscriptstyle T}})^{\mathrm{\scriptscriptstyle T}}\in\mathbb{R}^{n\times q}$ is a random effect design matrix.
Given the linear predictor $\bm{x}_{i}^{\mathrm{\scriptscriptstyle T}}\bm{\beta}+\bm{z}_{i}^{\mathrm{\scriptscriptstyle T}}\bm{u}$, the joint update of parameters related to dispersion ($\lambda$ or $(\bm{\lambda},\psi)$ and KG random variables is the same as fixed-effect models. Now, let the KG variables and dispersion parameters be fixed, then the likelihood (in terms of $\bm{\beta}$ and $\bm{u}$) is proportional to $\prod_{i=1}^{n}N(\tilde{y}_{i};\eta_{i},\kappa_{i}^{-1})$, where $\tilde{y}_{i}$ is defined as $\lambda(y_{i}-0.5)/\kappa_{i}$ for cobin and $\lambda_{i}(y_{i}-0.5)/\kappa_{i}$ for micobin. In matrix form, denoting $K=\hbox{diag}(\kappa_{1},\dots,\kappa_{n})$, the likelihood is proportional to $N_{n}(\tilde{\bm{y}};X\bm{\beta}+Z\bm{u},K^{-1})$.
Algorithm S.2 One cycle of a partially collapsed Gibbs sampler
1:Sample $\lambda$ from $\operatorname*{pr}(\lambda=l\mid\bm{\beta},\bm{u})\propto p_{\lambda}(l)\prod_{i=1}^{n}p_{\mathrm{cobin}}(y_{i};\eta_{i},l^{-1})$ among $\{1,\dots,L\}$
2:Sample $\kappa_{i}$ from $(\kappa_{i}\mid\lambda,\bm{\beta},\bm{u})\stackrel{{\scriptstyle\mathrm{ind}}}{{\sim}}\mathrm{KG}(\lambda,\eta_{i})$, $i=1,\dots,n$triangleright steps 1,2 jointly updates $(\lambda,\bm{\kappa})$
3:Sample $\bm{\beta}$ from $(\bm{\beta}\mid\lambda,\bm{\kappa},\vartheta)\sim N_{p}(\bm{m}_{\beta},V_{\beta})$, where $\triangleright$ $\bm{u}$ integrated out but conditioned on $\vartheta$
$V_{\beta}^{-1}=X^{\mathrm{\scriptscriptstyle T}}(K^{-1}+Z\Sigma_{u}(\vartheta)Z^{\mathrm{\scriptscriptstyle T}})^{-1}X+\Sigma_{\beta}^{-1},\quad\bm{m}_{\beta}=V_{\beta}X^{\mathrm{\scriptscriptstyle T}}(K^{-1}+Z\Sigma_{u}(\vartheta)Z^{\mathrm{\scriptscriptstyle T}})^{-1}\tilde{\bm{y}}$
4:Sample $\vartheta$ from $p(\vartheta\mid\bm{\beta},\lambda,\bm{\kappa})$ using Metropolis-Hastings with some proposal $\vartheta^{\star}\sim q(\vartheta^{\star}\mid\vartheta)$,
[…]
5:Sample $\bm{u}$ from $(\bm{u}\mid\vartheta,\bm{\beta},\lambda,\bm{\kappa})\sim N_{q}(\bm{m}_{u},V_{u})$, where $\triangleright$ Steps 4 and 5 jointly updates $\vartheta$ and $\bm{u}$
$V_{u}^{-1}=Z^{\mathrm{\scriptscriptstyle T}}KZ+\Sigma(\vartheta)^{-1},\quad\bm{m}_{u}=V_{u}Z^{\mathrm{\scriptscriptstyle T}}K(\tilde{\bm{y}}-X\bm{\beta})$
The sampler first updates $\bm{\beta}$ from the partially collapsed posterior $p(\bm{\beta}\mid\lambda,\bm{\kappa},\vartheta)$ where $\bm{u}$ is marginalized out but the random effect covariance parameter $\vartheta$ is not. By normal-normal conjugacy, integrating out $\bm{u}\sim N_{n}(\bm{0},\Sigma_{u}(\vartheta))$, we have
$p(\bm{\beta}\mid\lambda,\bm{\kappa},\vartheta)\propto N_{n}(\tilde{\bm{y}};X\bm{\beta},K^{-1}+Z\Sigma_{u}(\vartheta)Z^{\mathrm{\scriptscriptstyle T}})N_{p}(\bm{\beta};\bm{0},\Sigma_{\beta})$
which yields step 3 of Algorithm S.2. Next, we update $(\vartheta,\bm{u})$ jointly by first sampling $\vartheta\sim p(\vartheta\mid\bm{\beta},\lambda,\bm{\kappa})$ and then sampling $\bm{u}\sim p(\bm{u}\mid\vartheta,\bm{\beta},\lambda,\bm{\kappa})$. This is based on
$p(\bm{u},\vartheta\mid\bm{\beta},\lambda,\bm{\kappa})\propto N_{q}(\bm{u};(Z^{\mathrm{\scriptscriptstyle T}}K^{-1}Z)^{-1}Z^{\mathrm{\scriptscriptstyle T}}K(\tilde{\bm{y}}-X\bm{\beta}),(Z^{\mathrm{\scriptscriptstyle T}}K^{-1}Z)^{-1})N_{q}(\bm{u};\bm{0},\Sigma(\vartheta))p(\vartheta)$
where we used $N_{n}(Z\bm{u};\bm{\mu},K)\propto N_{q}\left(\bm{u};(Z^{\mathrm{\scriptscriptstyle T}}K^{-1}Z)^{-1}Z^{\mathrm{\scriptscriptstyle T}}K^{-1}\bm{\mu},(Z^{\mathrm{\scriptscriptstyle T}}K^{-1}Z)^{-1}\right)$ in terms of $\bm{u}$ for a full column rank $Z\in\mathbb{R}^{n\times q}$. Conditioning on $\vartheta$, it yields step 5 of Algorithm S.2. Marginalizing out $\bm{u}$ using normal-normal conjugacy, we have
$p(\vartheta\mid\bm{\beta},\lambda,\bm{\kappa})\propto\underbrace{N_{q}\left((Z^{\mathrm{\scriptscriptstyle T}}K^{-1}Z)^{-1}Z^{\mathrm{\scriptscriptstyle T}}K(\tilde{\bm{y}}-X\bm{\beta});\bm{0},(Z^{\mathrm{\scriptscriptstyle T}}K^{-1}Z)^{-1}+\Sigma(\vartheta)\right)}_{\mathcal{L}(\vartheta)}p(\vartheta),$ (S.16)
which gives step 4 of Algorithm S.2. While it is possible to condition on $\bm{u}$ when sampling $\vartheta$, which may provide a direct sampler (such as when $\vartheta$ is a marginal variance and with an inverse gamma prior), marginalizing out $\bm{u}$ significantly improves mixing.
Now we specialize this generic sampler into three cases.
Random intercept model. In this case, the dimension of the random effect coefficient $q$ ($q<n$) corresponds to the number of groups. Let $g_{i}\in\{1,\dots,q\}$ be a group label for $i=1,\dots,n$. Then the random effect design matrix $Z$ has elements $Z_{i,j}=1$ if $g_{i}=j$ and $Z_{i,j}=0$ otherwise. Importantly, $Z^{\mathrm{\scriptscriptstyle T}}K^{-1}Z=Z^{\mathrm{\scriptscriptstyle T}}\hbox{diag}(1/\kappa_{1},\dots,1/\kappa_{n})Z$ becomes a diagonal matrix.
Spatial mixed effects model. Spatial regression model (7) corresponds to the case when $q=n$ and $Z=I_{n}$. Here $Z^{\mathrm{\scriptscriptstyle T}}K^{-1}Z=K^{-1}$ is a diagonal matrix.
Spatial mixed effects model with sparse $\Sigma(\vartheta)^{-1}$. When the precision matrix $\Sigma(\vartheta)^{-1}$ is known to be sparse, such as under the NNGP model, matrix operations that involve inversion of $n\times n$ matrices can utilize sparse matrix algorithms. First, in step 2, the calculation of $V_{\beta}$ involves $(K^{-1}+\Sigma_{u}(\vartheta))^{-1}=K-K(\Sigma(\vartheta)^{-1}+K)^{-1}K$, where the equation follows from the Woodbury identity. Here $\Sigma(\vartheta)^{-1}+K$ is sparse since $K$ is diagonal, thus its inversion can utilize a sparse matrix algorithm. Next, in step 4, it involves the inverse and determinant calculation of the covariance matrix of the $\mathcal{L}(\vartheta)$ term, i.e. $(K+\Sigma(\vartheta))^{-1}=K^{-1}-K^{-1}(\Sigma(\vartheta)^{-1}+K^{-1})^{-1}K^{-1}$ and $|K+\Sigma(\vartheta)|=|\Sigma(\vartheta)^{-1}+K^{-1}||K||\Sigma(\vartheta)|$, which are all based on sparse matrices. Finally, in step 5, $K+\Sigma(\vartheta)^{-1}$ is already sparse. Leveraging such sparsity significantly improves computation, where we used spam R package (Furrer and Sain,, 2010). Note that this corresponds to a block update of $\bm{u}$ that is different from a sequential update of $\bm{u}$, a default sampler of spNNGP R package (Finley et al.,, 2022).
For a low-rank structure on a covariance matrix $\Sigma(\vartheta)$, a similar strategy involving Woodbury matrix identity can be employed; see Lee and Dunson, (2024) for details.
The derivations for micobin mixed effects models are essentially the same except for step 1 and an additional step of sampling $\psi$. Derivations for those additional steps are detailed in the previous subsection and thus omitted.
S.4.3 Varying dispersion micobin model
Next, we describe the posterior inference procedure for variable dispersion micobin regression (6). Given $\bm{\lambda}$, the dispersion submodel is a negative binomial regression model $(\lambda_{i}-1)\sim\mathrm{negbin}(2,\psi_{i})$, $\mathrm{logit}(\psi_{i})=\bm{d}_{i}^{\mathrm{\scriptscriptstyle T}}\bm{\gamma}$ Therefore, under the normal prior on $\bm{\gamma}\sim N(\bm{\mu}_{\gamma},\Sigma_{\gamma})$, we can combine Pólya-Gamma (PG) data augmentation (Polson et al.,, 2013) that leads to conditionally conjugate sampling of $\bm{\gamma}$. The resulting Gibbs sampler has 5 steps, where steps 1 to 3 are equal to Algorithm 2. Denote $\bm{d}_{i}\in\mathbb{R}^{d}$ as a covariate for the dispersion model, and $D\in\mathbb{R}^{n\times d}$ as a design matrix. The new step 4 corresponds to sampling $\omega_{i}\stackrel{{\scriptstyle\mathrm{ind}}}{{\sim}}\mathrm{PG}(1+\lambda_{i},\bm{d}_{i}^{\mathrm{\scriptscriptstyle T}}\bm{\gamma})$ for $i=1,\dots,n$. Step 5 corresponds to sampling $(\bm{\gamma}\mid\bm{\lambda},\bm{\omega})\sim N_{q}(\bm{m}_{\gamma},V_{\gamma})$, where $V_{\gamma}^{-1}=D^{\mathrm{\scriptscriptstyle T}}\hbox{diag}(\omega_{1},\dots,\omega_{n})D+\Sigma_{\gamma}^{-1}$ and $\bm{m}_{\gamma}=V_{\gamma}\{D^{\mathrm{\scriptscriptstyle T}}(1.5-0.5\lambda_{1},\dots,1.5-0.5\lambda_{n})^{\mathrm{\scriptscriptstyle T}}+\Sigma_{\gamma}^{-1}\bm{\mu}_{\gamma}\}$.
S.4.4 EM algorithm
The Kolmogorov-Gamma augmentation also leads to an EM algorithm that can be used to find the MLE or posterior mode under a cobin regression model with cobit link. First, when $\kappa\sim\mathrm{KG}(b,c)$, we have $E(\kappa)=bc^{-2}\{(c/2)\coth(c/2)-1\}$ if $c\neq 0$ or $E(\kappa)=b/12$ if $c=0$, which is easy to check from the definition of the KG as an infinite convolution of gammas.
Based on the prior $\bm{\beta}\sim N_{p}(\bm{0},\Sigma_{\beta})$ and the augmented model (11) with fixed $\lambda$, we are interested in finding the posterior mode
$\hat{\bm{\beta}}=\operatorname*{arg\,max}_{\bm{\beta}}\int_{\mathbb{R}_{+}^{n}}\exp(-\bm{\beta}^{\mathrm{\scriptscriptstyle T}}\Sigma_{\beta}^{-1}\bm{\beta}/2)\prod_{i=1}^{n}\left\{\exp(\lambda(y_{i}-0.5)\bm{x}_{i}^{\mathrm{\scriptscriptstyle T}}\bm{\beta}-\kappa_{i}(\bm{x}_{i}^{\mathrm{\scriptscriptstyle T}}\bm{\beta})^{2}/2){p_{\textsc{kg}}}(\kappa_{i};\lambda,0)\right\}\mathrm{d}\bm{\kappa}$
When finding MLE, the prior term $\exp(-\bm{\beta}^{\mathrm{\scriptscriptstyle T}}\Sigma_{\beta}^{-1}\bm{\beta}/2)$ is ignored. By Theorem 2, the conditional distribution of $\bm{\kappa}$ given $\bm{\beta}^{(t)}$ is $\kappa_{i}\stackrel{{\scriptstyle\mathrm{ind}}}{{\sim}}\mathrm{KG}(\lambda,\bm{x}_{i}^{\mathrm{\scriptscriptstyle T}}\bm{\beta}^{(t)})$ for $i=1,\dots,n$. Therefore, E step is
$Q(\bm{\beta}\mid\bm{\beta}^{(t)})=\mathrm{constant}-\frac{1}{2}\sum_{i=1}^{n}E(\kappa_{i})(\bm{x}_{i}^{\mathrm{\scriptscriptstyle T}}\bm{\beta})^{2}+\sum_{i=1}^{n}\lambda(y_{i}-0.5)\bm{x}_{i}^{\mathrm{\scriptscriptstyle T}}\bm{\beta}-\frac{1}{2}\bm{\beta}^{\mathrm{\scriptscriptstyle T}}\Sigma_{\beta}^{-1}\bm{\beta}$
Plugging in $\hat{\kappa}_{i}=\lambda(\bm{x}_{i}^{\mathrm{\scriptscriptstyle T}}\bm{\beta}^{(t)})^{-2}\{(\bm{x}_{i}^{\mathrm{\scriptscriptstyle T}}\bm{\beta}^{(t)}/2)\coth(\bm{x}_{i}^{\mathrm{\scriptscriptstyle T}}\bm{\beta}^{(t)}/2)-1\}$ in place of $E(\kappa_{i})$ (if $\bm{x}_{i}^{\mathrm{\scriptscriptstyle T}}\bm{\beta}^{(t)}=0$, $\hat{\kappa}_{i}=\lambda/12$), the M step is
$\bm{\beta}^{(t+1)}=(X^{\mathrm{\scriptscriptstyle T}}\hbox{diag}(\hat{\kappa}_{1},\dots,\hat{\kappa}_{n})X+\Sigma_{\beta}^{-1})^{-1}\{X^{\mathrm{\scriptscriptstyle T}}(\lambda\bm{y}-0.5\lambda\bm{1}_{n})\}$
Iterating the E step and M step (until $\bm{\beta}^{(t)}$ is stabilized given a tolerance level) gives the posterior mode (or MLE) of $\bm{\beta}$ under fixed $\lambda$. In fact, without the prior contribution $\exp(-\bm{\beta}^{\mathrm{\scriptscriptstyle T}}\Sigma_{\beta}^{-1}\bm{\beta}/2)$, the MLE does not depend on $\lambda$ due to orthogonality between $\bm{\beta}$ and $\lambda$.
For unknown $\lambda$ with candidates $\lambda\in\{1,\dots,L\}$ for some large $L$, one can run EM algorithm $L$ times, calculate the unnormalized posterior $p(\bm{\beta})\prod_{i=1}^{n}p_{\mathrm{cobin}}(y_{i};\bm{x}_{i}^{\mathrm{\scriptscriptstyle T}}\bm{\beta},\lambda)$, and choose $(\hat{\bm{\beta}}(\lambda),\lambda)$ that maximizes it.
S.5 Further details of spatial regression simulation study
We first describe the prior specification of dispersion parameters under three different regression models. First, for micobin regression, we consider the $\psi\sim\mathrm{Unif}(2,2)$ prior, reflecting the prior belief that $E\{\operatorname*{var}(Y)\}=0.5V(\mu)$ for a given mean $\mu=E(Y)$. For the prior of $\lambda$ for cobin regression, we consider $p(\lambda)=36\lambda\Gamma(\lambda+1)/\Gamma(\lambda+5)$ for $\lambda=1,2,\dots$, which is derived from the marginal distribution of $\lambda$ when $(\lambda-1)\mid\psi\sim\mathrm{Negbin}(2,\psi)$ and $\psi\sim\mathrm{Beta}(2,2)$. In practice, we truncate the support of $\lambda$ at some large upper bound $L$, where we choose $L=70$ throughout the paper.
The precision parameter $\phi$ in the beta regression model (the sum of two beta shape parameters) satisfies $\operatorname*{var}(Y)=\mu(1-\mu)/(1+\phi)$. While there are several different choices of $p(\phi)$ available, we choose squared uniform distribution $\phi\sim\mathrm{Unif}(0,A)^{2}$ for some $A>0$, following the suggestion of Figueroa-Zúñiga et al., (2013). To match the prior belief of $E(\psi)=1/2$ so that $E(\operatorname*{var}(Y))=0.5V(\mu)$, since $\mu(1-\mu)/3$ and $V(\mu)$ operates on a similar scale (see Figure S.5), we choose $A=8.74$ so that the prior $\phi\sim\mathrm{Unif}(0,8.74)^{2}$ satisfies $E(3/(1+\phi))\approx 1/2$. The beta regression model is fitted with Stan version 2.32.2 along with rstan version 2.32.6, with all default options using No-U-Turn sampler.
Next, we describe how negative test log-likelihood (negtestLL) and mean squared prediction error (MSPE) are calculated. Let $\{\bm{\theta}^{(m)}\}_{m=1}^{M}$ be a set of parameter samples from MCMC output. For each parameter sample $\bm{\theta}^{(m)}$, posterior predictive samples at new locations $\{s_{i}^{*}\}$ can be generated as $\eta(s_{i}^{*})^{(m)}=\bm{x}(s_{i}^{*})^{\mathrm{\scriptscriptstyle T}}\bm{\beta}^{(m)}+u(s_{i}^{*})^{(m)}$ for $i=1,\dots,n_{\mathrm{test}}$, where $(u(s_{1}^{*}),\dots,u(s_{n_{\mathrm{test}}}^{*}))^{(m)}$ is drawn from a multivariate normal conditioned on the $m$th sample of the spatial random effects at $n_{\mathrm{train}}$ locations. Based on the true mean $\mu(s^{*}_{i})$ and held-out realizations $y(s_{i}^{*})$ for $i=1,\dots,n_{\mathrm{test}}$, the negative test log-likelihood based on the assumed model $p(y(s_{i}^{*})\mid\bm{\theta}^{(m)})$ (beta, cobin, micobin) is calculated as
$\text{negtestLL}=\frac{1}{n_{\mathrm{test}}}\sum_{i=1}^{n_{\mathrm{test}}}\log\left\{\frac{1}{M}\sum_{m=1}^{M}p(y(s_{i}^{*})\mid\bm{\theta}^{(m)})\right\}.$
The MSPE is calculated as
$\text{MSPE}=\frac{1}{n_{\mathrm{test}}}\sum_{i=1}^{n_{\mathrm{test}}}\{\mu(s_{i}^{*})-\hat{\mu}(s_{i}^{*})\}^{2}$
where $\hat{\mu}(s_{i}^{*})$ is a posterior mean estimate of the conditional mean at a new location $s_{i}^{*}$, i.e. $M^{-1}\sum_{m=1}^{M}B^{\prime}(\eta(s_{i}^{*})^{(m)})$. The reported quantities in Table 2 are averaged over 100 replicates. Computations are carried out under the Intel(R) Xeon(R) Gold 6336Y 2.40GHz CPU environment.
S.6 Additional information for MMI data analysis
The benthic macroinvertebrate multivariate index (MMI) data from the 2017 National Lakes Assessment Survey (U.S. Environmental Protection Agency,, 2022) and lake watershed covariates from LakeCat data (Hill et al.,, 2018) are joined based on the unique identifier of the lake (COMID). In Table S.2, we present three lakes that exhibit zero or very low MMI values, while Table S.3 outlines the details of the 9 selected LakeCat covariates.
Table S.2: Lakes with zero or very low MMI values (multiplied by 0.01 from the original scale). COMID: lake unique identifier used in NLA 2017 and LakeCat; Site ID: lake unique ID used in NLA;
COMID Site ID MMI longitude latitude Lake name County State
22721231 NLA17_GA-10002 0 -81.893 33.300 Brierpatch lake Richmond GA
9201925 NLA17_NC-10015 0.02 -80.263 35.116 Jones pond Anson NC
22845861 NLA17_AR-10001 0.021 -92.273 34.538 Ferguson lake Saline AR
Table S.3: Description of covariates from LakeCat data (Hill et al.,, 2018). NLCD refers to the National Land Cover Dataset. Watershed refers to a set of hydrologically aggregated catchments (including lake waterbody surface area) that represent the full contributing landscape area to a downslope lake. All variables are $\log_{2}(x+1)$ transformed.
Variable (unit) Brief and detailed description
agkffact (unitless) Ag soil erodibility Kf factor Mean of state soil geographic Kf factor raster on agricultural land (NLCD 2006) within the watershed. Kf factor is a soil erodibility factor which quantifies the susceptibility of soil particles to detachment and movement by water. This factor is used in the universal soil loss equation to calculate soil loss by water
bfi (%) Base flow index The component of streamflow that can be attributed to ground-water discharge into streams. bfi is the ratio of base flow to total flow within the watershed
[…]
conif (%) Evergreen forest percentage % of watershed area classified as evergreen forest land cover (NLCD class 42) in 2016
crophay (%) Row crop and pasture/hay percentage % of watershed area classified as crop and hay land use (NLCD classes 81, 82) in 2016
[…]
pestic97 (kg/km2) Mean pesticide use Mean pesticide use in year 1997 within the watershed
urbmdhi (%) Developed, medium and high intensity land use percentage % of watershed area classified as developed, medium and high intensity land use (NLCD classes 23, 24) in 2016
We standardized all (log-transformed) covariates to have a mean of 0 and a variance of 1 prior to running MCMC, and transformed back to the original scale after sampling. Based on these standardized covariates, we used the same prior as in the simulation study. That is, $\bm{\beta}\sim N_{p}(\bm{0},100^{2}I_{p})$ for regression coefficients, a standard half-Cauchy prior on the spatial random effect standard deviation $\sigma_{u}$, beta dispersion $\phi\sim\mathrm{Unif}(0,8.74)^{2}$, cobin dispersion $p(\lambda)\propto\lambda\Gamma(\lambda+1)/\Gamma(\lambda+5)$ for $\lambda=1,2,\dots$, and micobin dispersion $\psi\sim\mathrm{Beta}(2,2)$. We set the upper bound of $\lambda$ to be $L=70$.
The MMI data analysis involves an NNGP prior for the spatial random effects. To enable nearest neighbor calculation using the spNNGP R package (Finley et al.,, 2022), we used WGS 84 / UTM zone 15N (EPSG:32615) coordinate system. The NNGP prior requires the specification of an ordering and a number of neighborhoods. Following the default setting of spNNGP R package, we use coordinate-based ordering and use 15 nearest neighbors. We use an exponential covariance kernel $\operatorname*{cov}\{u(s),u(s^{\prime})\}=\sigma_{u}^{2}\exp(-\|s-s^{\prime}\|/\rho)$ in the construction of the NNGP prior. We fixed the spatial range parameter at $\rho=200$km, leading to an effective range of approximately 600km, the distance where the spatial correlation is below 0.05 (before NNGP approximation). This specification is consistent with the analysis of Fox et al., (2020), where the empirical spatial autocorrelation of stream benthic macroinvertebrate MMI in the US was found to be close to zero for distances beyond 580km.
In the comparative analysis shown in Table 3 in the main manuscript, the lake with a zero MMI (COMID: 22721231) was excluded since beta and cobin regressions do not accommodate an exact 0 response. Table S.4 summarizes the spatial micobin regression results using the full dataset ($n=950$), along with results from Table 3 ($n=949$) for a comparison. The estimates and credible intervals are highly similar to each other. Furthermore, the quantile residual plot in Figure 4 indicates that two low-MI lakes (COMID: 9201925, 22845861) are highly influential in the beta and cobin regression fit with $n=949$. After removing these two lakes, the beta, cobin, micobin regression results based on the data with $n=947$ are presented in Table S.5.
For the beta regression results, the selected variables based on 95% credible intervals not including zero changed; bfi is no longer selected, while agkffact and conif are now selected. For the cobin and micobin regressions, there were no changes in selected variables. In terms of the difference in the posterior mean estimate $\hat{\bm{\beta}}$ compared to the result with data of size $n=949$, measured by the Euclidean distance, the beta regression exhibits the largest change, with a value of 0.743 (0.516 excluding the intercept). The cobin regression shows a change of 0.444 (0.273 excluding the intercept), and the micobin regression is the most stable, with a change of 0.069 (0.040 excluding the intercept).
Table S.4: MMI data analysis results with original data with size $n=950$ using micobin regression, along with the result from Table 3 ($n=949$) for a comparison. Results where 95% credible intervals that do not include zero are highlighted in bold.
Micobin regression $(n=950)$ Micobin regression $(n=949)$
Variable Estimate 95% CI Estimate 95% CI
(Intercept) -1.758 (-3.517, -0.04) -1.797 (-3.551, -0.085)
agkffact -3.456 (-6.175, -0.794) -3.457 (-6.113, -0.800)
bfi 0.219 (-0.100, 0.537) 0.229 (-0.082, 0.548)
cbnf 0.187 (-0.040, 0.415) 0.191 (-0.035, 0.425)
conif 0.128 (0.048, 0.208) 0.123 (0.044, 0.203)
crophay -0.060 (-0.222, 0.101) -0.054 (-0.213, 0.105)
fert -0.071 (-0.296, 0.13) -0.082 (-0.300, 0.138)
manure -0.031 (-0.178, 0.115) -0.029 (-0.173, 0.118)
pestic97 -0.023 (-0.106, 0.059) -0.025 (-0.108, 0.059)
urbmdhi -0.141 (-0.243, -0.038) -0.142 (-0.243, -0.041)
agkffact, soil erodibility factor; bfi, base flow index; cbnf, cultivated biological N fixation; conif, coniferous forest cover; crophay, crop/hay land cover; fert, synthetic N fertilizer use; manure, manure application; pestic97, 1997 pesticide use; urbmdhi, medium/high-density urban land cover. All variables are $\log_{2}(x+1)$ transformed.
Table S.5: Comparison of MMI data analysis results with three different models with dataset of size $n=947$. Variables where 95% credible intervals that do not include zero are highlighted in bold. The last row corresponds to the Euclidean distance between the estimates based on $n=949$ and $n=947$.
$(n=947)$ Beta regression Cobin regression Micobin regression
Variable Estimate 95% CI Estimate 95% CI Estimate 95% CI
(Intercept) -1.829 (-3.621, -0.070) -1.756 (-3.531, 0.006) -1.741 (-3.520, 0.018)
agkffact -3.088 (-5.982, -0.206) -3.150 (-6.019, -0.258) -3.494 (-6.220, -0.822)
bfi 0.244 (-0.075, 0.568) 0.228 (-0.088, 0.552) 0.219 (-0.097, 0.540)
cbnf 0.191 (-0.044, 0.430) 0.200 (-0.035, 0.437) 0.196 (–0.034, 0.424)
conif 0.096 (0.014, 0.175) 0.103 (0.021, 0.183) 0.125 (0.045, 0.204)
crophay -0.057 (-0.223, 0.110) -0.053 (-0.218, 0.114) -0.050 (-0.210, 0.110)
fert -0.096 (-0.327, 0.135) -0.104 (-0.329, 0.122) -0.089 (-0.316, 0.135)
manure -0.001 (-0.148, 0.148) -0.009 (-0.157, 0.138) -0.022 (-0.167, 0.122)
pestic97 -0.031 (-0.118, 0.057) -0.030 (-0.119, 0.057) -0.027 (-0.110, 0.055)
urbmdhi -0.180 (-0.283, -0.076) -0.169 (-0.275, -0.064) -0.143 (-0.242, -0.043)
Change $\|\hat{\bm{\beta}}^{(949)}-\hat{\bm{\beta}}^{(947)}\|_{2}=0.743$ $\|\hat{\bm{\beta}}^{(949)}-\hat{\bm{\beta}}^{(947)}\|_{2}=0.444$ $\|\hat{\bm{\beta}}^{(949)}-\hat{\bm{\beta}}^{(947)}\|_{2}=0.069$
agkffact, soil erodibility factor; bfi, base flow index; cbnf, cultivated biological N fixation; conif, coniferous forest cover; crophay, crop/hay land cover; fert, synthetic N fertilizer use; manure, manure application; pestic97, 1997 pesticide use; urbmdhi, medium/high-density urban land cover. All variables are $\log_{2}(x+1)$ transformed.